### v1

# Towards Internet-Scale Training For Agents

Abstract
The predominant approach for training web navigation agents gathers human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data are an inefficient resource. We develop a pipeline to facilitate Internet-scale training for agents without laborious human annotations. In the first stage, an LLM generates tasks for 150k diverse websites. In the next stage, LLM agents complete tasks and produce trajectories. In the final stage, an LLM reviews the trajectories and judges their success. Language models are competitive with human annotators, detecting and filtering out harmful content with an accuracy of 97%, generating feasible tasks with an 89% rate, and judging successful trajectories with an 82.6% accuracy. Scaling the pipeline, agents based on Llama 3.1 70B solve 16.7% of tasks for 150k sites. Training on the data generated by our pipeline is competitive with training on human demonstrations. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data. When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web. Code will be available at: data-for-agents.github.io.
1 Introduction
The predominant approach to training LLM-based web navigation agents is to collect human demonstrations on a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023). Human data can be laborious to collect and becomes costly to scale as the breadth of skills that users require from language model agents grows. There are more than three-hundred million sites on the Internet (The Common Crawl Foundation, 2024), and the range of sites that researchers can manually prepare for human annotation represents a tiny fraction of the Internet. The key problem is that human data can become unreliable at scale. Human-written web navigation tasks are highly effective for popular sites, but reliability drops for sites with lower popularity due to annotators’ lack of familiarity. For these sites with lower popularity, which represent the majority of sites on the internet, human-written web navigation tasks are feasible only 40% of the time (discussed in Section 4), requiring a costly manual verification step. For the same pool of sites, language models improve feasibility rates to more than 80%. There is a growing need to automate data pipelines for the next generation of agents trained at internet scale. We address a key challenge by reducing the dependency on human data in the agent pipeline. We develop an automatic data pipeline that aims to facilitate Internet-Scale Training for Agents, which we shorten to InSTA, a pipeline that relies on synthetic web navigation tasks proposed, attempted, and evaluated by language models.
Figure 1: Overview of the proposed agent pipeline. We develop a pipeline for training web navigation agents at internet scale using tasks proposed, attempted, and evaluated by pretrained large language models. We generate 150k diverse tasks across 1M internet sites. Code for our data generation pipeline, and traces for agent rollouts will be available on our website: data-for-agents.github.io.
Our method operates in three stages. In the first stage, we employ a language model to propose candidate web navigation tasks for an agent to perform across 150k live sites on the Internet. Current works are limited to 200 popular sites (Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023) that human annotators are likely to be familiar with. Language models help us scale to 1,000 times more sites than current efforts, with better coverage of real-world sites. One major consideration when scaling up training for agents is safety: building safe agents requires that we avoid sites with harmful, unsafe, or dangerous content. We evaluated the ability of language models to detect such content and aggressively filtered out 85% candidates from the initial 1M sites to 150k sites that are judged safe by language models. These models succeed at detecting safe content with an accuracy of 97%, compared to 75% human accuracy. With tasks generated across a safe and diverse set of websites, we proceed to run language model agents to attempt the generated tasks.
In the second stage of the pipeline, a language model agent attempts to complete tasks using a web browser. We provide the entire Playwright API to the agent, which operates the browser by generating function calls in the Playwright API, a standard choice in recent web navigation benchmarks (Chezelles et al., 2024; Du et al., 2024; Drouin et al., 2024). While existing LLM agents can be directly applied to our tasks, measuring their success presents another challenge: sites organize data in bespoke formats. Previous efforts circumvent this by enlisting human labelers to review attempts manually and judge their success (Lù et al., 2024; Deng et al., 2023), or by designing sites for benchmarking where formats are consistent (Zhou et al., 2024b; Koh et al., 2024a). However, neither approach scales efficiently, and human annotators may not be reliable for sites they are not familiar with in the tail of the distribution. In the third stage of the pipeline, we scale the evaluation using language models. We employ LLMs to judge (Lightman et al., 2024) whether a task is solved by the final timestep and obtain an accuracy of up to 93.1% in detecting successful trajectories for the most confident predictions. Llama-3.1-70B-Instruct solves 16.7% of zero-shot tasks with a judge confidence of conf = 1. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1%, respectively, when training agents on mixtures of data from our pipeline and human data. When training agents with all human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
Our work develops InSTA, an automated data pipeline that aims to facilitate Internet-scale training for agents. Driven by pretrained language models, we generate web navigation tasks for 150k live websites, run LLM agents, and judge their success. By removing humans from the agent pipeline, we can scale and extend coverage towards sites where human data is less reliable. We present an analysis to improve safety and show that our data can outperform human demonstrations for training agents. Code for the pipeline will be available at: data-for-agents.github.io.
2 Related Works
Language Model Agents.
There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b) as the backbone for agents (Andreas, 2022). These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where text formats differ significantly from their training data. Search algorithms provide a secondary axis to improve the reasoning capabilities of language model agents (Yao et al., 2023b; Besta et al., 2024; Koh et al., 2024b; Zhou et al., 2024a) by providing an explicit algorithmic scaffold and allowing test-time compute to improve reasoning steps (Snell et al., 2024; Zhong et al., 2024). Although most of the work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024) as target benchmarks are becoming more difficult for zero-shot language models.
Agent Pipelines.
There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024). However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023), or the same tasks (Putta et al., 2024). We instead consider a setting where tasks and environment configurations are entirely separate between training and testing, creating a strong train-test split that follows recommended practice. This presents a challenge: the web navigation data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024). We address this challenge by increasing the scale of data collection and better coverage of the distribution of real-world sites. We train on diverse tasks generated by our pipeline and successfully transfer agents trained on our data to downstream benchmarks while maintaining a strong train-test split. Our training procedure resembles a modified FireAct (Chen et al., 2023), where language models jointly propose and evaluate tasks for agents.
Agent Datasets.
The majority of datasets for training web navigation agents rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024). This approach has limits, as the breadth and diversity of tasks researchers can manually curate are dwarfed by the volume of sites on the Internet. There are more than three-hundred million sites on the Internet according to The Common Crawl Foundation (2024), and existing datasets are limited to about 150 popular sites that human annotators are already familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024). There are hypothetical 1,000,000 times more data that could be available if we can efficiently harness this previously untapped resource. However, most sites are relatively obscure and human annotators are unreliable for sites they are not already familiar with. Finding suitable annotators becomes impractical at this massive scale, so we adapt language models to propose, attempt, and evaluate web navigation tasks. Although we are not the first to consider synthetic data for training agents (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), we have developed a key approach to harness internet-scale data efficiently.
Language Model Judges.
Core to our pipeline is a language model evaluator. Using language models to judge the correctness of responses is becoming popular to improve accuracy for LLMs (Li et al., 2024), and applications include verifying reasoning steps (Zhang et al., 2024), rejection sampling (Snell et al., 2024; Sun et al., 2024), prioritizing frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filtering out harmful responses (Inan et al., 2023), providing feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and providing ratings for alignment (Lee et al., 2024; Ouyang et al., 2024). Our use of language models to evaluate agent tasks is inspired by the generative verifier in Zhang et al. (2024), and modified from the multimodal verifier in He et al. (2024), where our language model predicts a confidence score that a task is solved, which is used to identify successful attempts.
3 Language Model Agents
Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a web environment in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$. Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{sys}}$.
$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}\big{(}\;\text{LLM}([\;\mathbf{x}_{\text{sys}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;])\;\big{)}$ (1)
Environment representations for observations and actions typically differ from the expected input format of the language model (typically images and text), and functions are introduced that map the observations to a multimodal prompt $\text{Enc}(\cdot)$, and parse actions from the language model generated response $f^{\;\text{text}\to\text{act}}(\cdot)$. For web navigation, the environment state $\mathbf{s}_{t}$ is HTML DOM, and is often formatted as raw HTML code, an Accessibility Tree, Set-of-marks, or screenshots (Zhou et al., 2024b; Koh et al., 2024a; Chezelles et al., 2024; Shen et al., 2024). In this work, we employ an off-the-shelf HTML-to-Markdown parser to convert observations into a textual format. Action formats vary between works, and we build on Schick et al. (2023)’s function-calling framework, where a language model generates code that is parsed into a function name and corresponding arguments. Given a set of strings $L$, and a set of function argument values $G$, the set of actions $\mathcal{A}$ is:
$\mathcal{A}=\big{(}L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\big{)}$ (2)
where $L_{\text{func}}$ is the set of function names in our API, and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$). We provide agent access to the entire Playwright API (Microsoft, 2024), a Microsoft-developed browser automation library that wraps around a Chromium web browser. The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the task is complete, after which point the agent calls stop with an optional answer:
[…]
We prompt the language model backbone to generate actions wrapped in a JSON code block for straightforward parsing. The action parser $f^{\;\text{text}\to\text{act}}$ consists of a regex template that matches the first JSON code block, such as the example in Figure 1, followed by JSON decoding of the string contents within the code block. When parsing fails due to invalid syntax, we generate a new response until parsing succeeds. Equipped with a language model agent that makes calls to the Playwright API, we face a crucial roadblock that impedes scaling: obtaining large and diverse data.
4 Internet-Scale Task Generation
Building internet-scale agents requires a diverse scaffold of tasks and environment configurations beyond what can be attained via manually curated examples annotated by humans. We develop a pipeline to efficiently harness vast quantities of sites on the Internet that aims to facilitate Internet-Scale Training for Agents, shortened to InSTA. Our pipeline uses pretrained language models to generate, attempt, and evaluate synthetic web navigation tasks for a more diverse pool of sites than current efforts that rely on tasks designed by researchers (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024). Human data is a valuable resource, but can be inefficient to gather. Our work shows that language models can be as accurate as human annotators. By removing human data from the agent pipeline, we can improve the safety and reliability of tasks, and efficiently scale task generation to 1M sites.
4.1 Language Model Task Proposer
In the first stage, we scale the generation of web navigation tasks to 1M diverse websites using a Language Model Task Proposer. Shown in Figure 2, the task proposer serves two key functions in the pipeline: (1) filtering out harmful content and websites that cannot be safely annotated, and (2) proposing realistic web navigation tasks that a hypothetical user might want to accomplish.
Figure 2: Task proposal and filtering for 150k live websites. Starting from 1,000,000 websites, we employ a pretrained language model that marks sites as safe/unsafe for annotation, and assigns a realistic task that a hypothetical user might want to accomplish on each site. The task proposer aggressively filters out 85% of websites from the pipeline, resulting in 150k safe websites annotated with realistic tasks.
Model Details.
We utilize pretrained and frozen language models that conform to a chat interface and accept a system prompt $\mathbf{x}_{\text{sys}}$, and a series of in-context examples through interleaved user and assistant prompts $\mathbf{x}_{\text{usr}}$ and $\mathbf{x}_{\text{ast}}$. The system prompt used for task generation is shown in Figure 2, and outlines all cases for which sites are considered unsafe for annotation. We consider the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b; a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google. Inference is served using vLLM (Kwon et al., 2023) for the Llama series of models. We employ a sampling temperature of 0.5, and a maximum budget of 64 newly generated tokens; all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
Prompt Details.
The goal of the task proposer is to accurately detect unsafe websites and generate realistic web navigation tasks when appropriate. We prompt the task proposer with the system prompt in Figure 2, a series of in-context examples (listed in Appendix E), and a final user prompt containing just the URL of the target website. We instruct the LLM via the system prompt to provide a task for the target website or to return “N/A” and mark the website as not suitable for annotation. This format produces a throughput of 20 websites per second for Llama 3.1 70B served on 16 GPUs with vLLM, processing 1M sites in 14 hours. The efficiency of stage one aids in scaling to large numbers of sites on the Internet, but we must not compromise safety and reliability for efficiency. To understand the trade-offs presented by our task proposal approach, we compare against typical human annotators at detecting safe websites for annotation and creating realistic agent tasks.
Method Acc. Prec. Recall
Llama 3.1 70B 85% 0.77 1.00
GPT-4o 95% 0.91 1.00
Gemini 1.5 Pro 97% 0.96 0.98
Human Baseline 75% 0.71 0.84
Figure 3: Accuracy for detecting harmful sites. We select 100 website domains, where 50 are safe, and 50 are unsafe based on the criteria in Figure 2. Pretrained language models exceed the accuracy and recall of human annotators at detecting harmful sites that are unsuitable for training agents safely.
4.2 Improving Safety
Language models beat single pass non-expert annotators at detecting websites suitable for annotation. To evaluate detection performance, we employ the task proposer as a classifier and consider sites where the task proposer returns “N/A” as the positive class. We curate 50 safe and 50 unsafe domains, based on the filtering conditions outlined in the system prompt in Figure 2 (selected websites and their URLs are listed in Appendix E). We generate task proposals for each site and measure the accuracy, precision, and recall of our safety filter compared to human annotators. The annotators are asked to classify each site as suitable or unsuitable for annotation based on the website URL, and the criteria listed in the system prompt the same observations given to the task proposer to ensure a fair comparison. Results are presented in Table 3.
Understanding The Results.
Language models outperform human annotators by 29.3% in accuracy, 35.2% in precision, and 31.0% in recall at detecting harmful sites. While larger models like Gemini 1.5 Pro show the best overall accuracy, smaller models like Llama 3.1 70B display high recall with a minor decrease in accuracy. Recall matters most for safety filters, and these results suggest that Llama 3.1 70B is sufficient to detect most harmful sites with high confidence.
4.3 Improving Reliability
Language models are more reliable than human annotators at creating realistic web navigation tasks. To evaluate reliability, we measure the rate at which human workers are able to perform web navigation tasks generated by our pipeline. We curate a set of 100 safe website domains (different from the safety experiment, refer to Appendix E), generate task proposals using our pipeline, and measure the rate of self-reported task completion for human workers performing tasks. Workers start from the initial website URL in their browser and navigate pages using their mouse and keyboard while staying on the original site, reporting once the task is complete or once they believe the task is not feasible on the current site. We compare the feasibility rates for tasks generated by our pipeline and tasks written by human annotators given the criteria listed in Figure 2. Results are shown in Table 4.
Method Feasibility Rate
Llama 3.1 70B 75%
GPT-4o 85%
Gemini 1.5 Pro 89%
Human Baseline 54%
Figure 4: Expert feasibility of proposed tasks. We generate tasks for 100 safe websites (listed in Appendix E), and measure the completion rates of human workers attempting to perform the generated tasks in their browser. Language models exceed the performance of human annotators at creating feasible web navigation tasks for agents.
Understanding The Results.
Language models outperform human annotators by 64.8% at creating feasible web navigation tasks. Larger models like Gemini 1.5 Pro display the best feasibility rates, but the smaller model Llama 3.1 70B still outperforms the human annotators by 38.9%. To understand the relationship between the popularity of the site being annotated and the reliability of human-written tasks, we conduct an experiment in Figure 5 comparing PageRank values (Page et al., 1999) of sites according to the official June 2024 host-level web graph from The Common Crawl Foundation (2024), versus the feasibility rates of the proposed tasks from Table 4.
Although human annotators match the reliability of LLMs at creating feasible web navigation tasks for popular sites, LLMs outperform human annotators by 157.1% for less popular sites with low PageRank values. As obscurity increases, human annotators are less familiar with the sites, and the reliability of their task proposals decreases by 55.7%, whereas the reliability of tasks generated by LLMs remains relatively constant. This difference suggests that we should employ language models to ensure that task proposals are reliable as we begin to scale agents to vast numbers of sites on the Internet. However, where do we acquire this large and diverse set of websites to process for annotation?
Figure 5: Feasibility rates vs PageRank values. We visualize PageRank values, a proxy for the popularity of websites, versus the expert feasibility rates of proposed web tasks. Human-written tasks perform on par with LLMs for popular sites, but as target sites become less popular and annotators are less familiar with them, LLMs begin to outperform human annotators at creating feasible tasks for agents.
4.4 Scaling To 150,000 Sites
We propose to leverage open-source crawls of the internet for large-scale task generation. As of June, 2024, the web graph released by The Common Crawl Foundation (2024) contains more than 300 million unique hosts, which we adapt to a data source for agents. In particular, we sort hosts by their PageRank values, and select the top 1M sites for task generation. CommonCrawl is likely to contain many sites not suitable for annotation, and the experiments in Section 4.2 illustrate that the safety filter in the task proposer can detect and remove them effectively. In our configuration, task generation with Llama 3.1 70B takes 14 hours for 1M sites served with vLLM (Kwon et al., 2023) on two 8-GPU nodes. Sections 4.2 and 4.3 show Llama 3.1 70B outperforms human annotators in safety and reliability, and we can serve it locally at significantly reduced cost compared to proprietary LLMs with a marginal loss in quality. The distribution of tasks generated with Llama 3.1 70B for the top 1M sites in the CommonCrawl PageRank are visualized in Figure 6.
Understanding The Data.
The task proposer filters out 85% of the sites in CommonCrawl, resulting in 150k sites that can be safely assigned tasks for agents. Visualized in Figure 6, our distribution is denser than human-written tasks, has a wide coverage of real-world sites, and diverse categories of tasks. We automatically label task categories (procedure in Appendix E) and find that 89% of categories have fewer than the mean of 16.9 tasks per category. Top categories include news search, recipe search, product lookup, tutorial search, event schedules, health information, and many more. Refer to Appendix E for the top categories. Empowered by this large and diverse collection of tasks across the internet, we can start building internet-scale agents.
Figure 6: Distribution of 150k tasks. We compare the distribution of tasks generated by our pipeline (blue points) to the Mind2Web (Deng et al., 2023) dataset (orange points) via textual features extracted by a sentence embedding model, and projected in 2D with UMAP (McInnes et al., 2020). Our distribution is denser than human-written tasks, and has broad coverage of diverse real-world sites and tasks.
5 Internet-Scale Agents
In the next stage, we begin to scale LLM agents to diverse web navigation tasks across the web. Shown in Figure 7, we initialize a web browsing environment to the URL provided to the task proposer in Section 4, and run a language model agent to complete tasks by generating function calls in the Playwright API. For evaluation, current efforts typically use human-written constraints based on the final URL or page state (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; Drouin et al., 2024), but it can be difficult to scale these. Recall from Figure 5 that human annotators are less reliable for sites lower in the PageRank, where their familiarity is reduced. Results in section 4 showed that language models beat humans in safety and reliability for task generation. As we begin to scale agents to diverse websites from the internet, can we replace human-written criteria with language model judgments for efficient evaluation? Their robustness remains an important unresolved question, and current work only considers language model judges for a limited set of popular websites (He et al., 2024). We begin by validating the robustness of language models for evaluating diverse internet tasks.
Figure 7: Automatic evaluation for agents with language model judges. Building on the large and diverse set of tasks generated by the pipeline, we employ pretrained language models to attempt and evaluate web navigation tasks. We dispatch language model agents to perform tasks by making calls to the Playwright API. We then employ language model judges to evaluate the trajectories.
5.1 Evaluation With Language Models
We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved by the final timestep. Conditioned on a system prompt $\mathbf{y}_{\text{sys}}$, a task $\mathbf{c}$, the history of previous actions $\mathbf{a}_{1},\cdots,\mathbf{a}_{T}$ and the final observation from a web browsing environment $\mathbf{s}_{T}$ formatted in text, we generate a completion using a language model and employ a function $f^{\;\text{text}\to\text{val}}(\cdot)$ to parse the estimated probability.
$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}\big{(}\;\text{LLM}([\;\mathbf{y}_{\text{sys}},\mathbf{c},\mathbf{a}_{1},\cdots,\mathbf{a}_{T},\text{Enc}(\mathbf{s}_{T})\;])\;\big{)}$ (4)
Building on the sites used to measure reliability in Section 4.3, we conduct an experiment to measure the accuracy of language models to detect successful web navigation trajectories. We run agents on tasks generated by Llama 3.1 70B for the 100 sites in Section 4.3, and prompt language models to estimate the probability that tasks are solved $\mathbf{r}_{T}$ based on Equation 4. We then conduct human evaluations for the trajectories and manually assign binary success labels. Accuracy is calculated by applying a threshold to the predictions $\mathbf{r}_{T}>0.5$ to assign classes and tracking the rate at which the predictions agree with human labels. To understand robustness for sites of varying popularity, we report the accuracy of language models versus the PageRank of the corresponding sites. Similarly, to understand the ability of language models to judge their own uncertainty, we report their accuracy versus their prediction confidence, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$ (twice the total variation distance from the uniform distribution to the predicted distribution). Results are shown in Figure 8.
Figure 8: Language models are robust evaluators. We measure the accuracy of language models for detecting successful trajectories, and find that accuracy remains stable relative to PageRank values (left plot). As models become more confident, their accuracy improves (right plot), suggesting confidence is a useful proxy for the reliability of their predictions.
Understanding The Results.
Language models are robust evaluators for web navigation tasks. Accuracy remains stable relative to PageRank values, suggesting that language models are effective for sites that typical human annotators are less familiar with. The best results are obtained with an evaluator based on GPT-4o, which attains an accuracy of $82.6\%$, compared to $81.7\%$ for Llama 3.1 70B, and $78.0\%$ for Gemini 1.5 Pro. Although the accuracy is robust to PageRank, the accuracy is highly informed by confidence. Language models show improved accuracy as their confidence improves, suggesting that they can effectively determine when their predictions are reliable. When considering predictions with conf = 1, the Llama 3.1 70B evaluator displays a compelling $93.1\%$ accuracy, 0.87 precision, and 0.82 recall for detecting successful trajectories. Now that we can efficiently and accurately judge trajectories, we can begin to scale language model agents to diverse internet tasks and track their success. Harnessing this judge, we can study the current abilities and shortcomings of language model agents spanning 150k websites.
Figure 9: Scaling LLM agents to 150k live sites. We run agents based on Llama 3.1 70B to complete tasks generated by our pipeline. We estimate success probabilities using a language model evaluator (left plot), and estimate probabilities agents are on the right track (right plot). 16.7% of tasks are estimated to be successful with conf = 1, and the spread of probabilities suggests data spans many difficulties.
5.2 Scaling To 150,000 Agents
We scale language model agents to 150k live sites in diverse domains across the Internet and attempt to complete 150k web navigation tasks generated by our pipeline. Shown in Figure 9, we evaluate the trajectories using a Llama 3.1 70B judge and run agents based on Llama 3.1 70B, selected because this model demonstrates high accuracy in Figure 8, and running currently available propriety models would be prohibitively expensive at this scale (see Appendix K for a cost analysis with different LLMs). We find that agents solve 16.7% of tasks with a model confidence of conf = 1. Furthermore, we observe that 35k tasks are judged to be on the right track with a confidence of conf = 1, suggesting that these could be solved if a larger computing budget was allocated. The spread along the x-axis in both plots in Figure 9 suggests that our tasks cover a broad range of difficulties, and working to solve them presents an opportunity to improve the capabilities of LLM agents. We observe that, when judging success, our evaluator tends to prefer binary predictions with high confidence values, suggesting that this subset of predictions is accurate based on Figure 8 results. Analyses for the agents that produced Figure 9 are presented in Appendix F.
Assembling the stages of the InSTA pipeline, we first generated tasks for 1M diverse sites across the Internet, we then dispatched language model agents to complete tasks for 150k sites marked as safe, and finally we evaluated the trajectories using a language model judge. Equipped with this large and diverse source of data, we seek to understand the utility of the data for training agents.
6 Training Agents
We compare the performance of agents trained on data from the InSTA pipeline to agents trained on human demonstrations from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), two recent and popular benchmarks for web navigation. Recent works that mix synthetic data with real data control the real data sampling probability in the batch $p_{\text{real}}$ independently from data size (Trabucco et al., 2024). We employ $p_{\text{real}}=0.5$ in few-shot experiments and $p_{\text{real}}=0.8$ otherwise. Shown in Figure 9, our data have a wide spread in performance, so we apply several filtering rules to select high-quality training data. First, we require the evaluator to return conf = 1 that the task was successfully completed, and that the agent was on the right track (this selects data where the actions are reliable, and directly caused the task to be solved). Second, we filter data where the trajectory contains at least three actions. Third, we remove data where the agent encountered any type of server error, was presented with a captcha, or was blocked at any point. These steps produce $7,463$ high-quality demonstrations in which agents successfully completed tasks on diverse websites. We sample 500 demonstrations uniformly at random from this pool to create a diverse test set, and employ the remaining $6,963$ demonstrations to train agents on a mix of real and synthetic data.
6.1 Improving Data-Efficiency
Figure 10: Data from InSTA improves efficiency. Language model agents trained on mixtures of our data and human demonstrations scale faster than agents trained on human data. In a setting with 32 human actions, adding our data improves Step Accuracy by +89.5% relative to human data for Mind2Web, and +122.1% relative to human data for WebLINX.
In a data-limited setting derived from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), agents trained on our data scale faster with increasing data size than human data alone. Without requiring laborious human annotations, the data produced by our pipeline leads to improvements on Mind2Web that range from +89.5% in Step Accuracy (the rate at which the correct element is selected and the correct action is performed on that element) with 32 human actions, to +77.5% with 64 human actions, +13.8% with 128 human actions, and +12.1% with 256 human actions. For WebLINX, our data improves by +122.1% with 32 human actions, and +24.6% with 64 human actions, and +6.2% for 128 human actions. Adding our data is comparable in performance gained to doubling the amount of human data from 32 to 64 actions. Performance on the original test sets for Mind2Web and WebLINX appears to saturate as the amount of human data increases, but these benchmark only test agent capabilities for a limited set of 150 popular sites.
6.2 Improving Generalization
Figure 11: Our data improves generalization. We train agents with all human data from the WebLINX and Mind2Web training sets, and resulting agents struggle to generalize to more diverse test data. Adding our data improves generalization by +149.0% for WebLINX, and +156.3% for Mind2Web.
To understand how agents trained on data from our pipeline generalize to diverse real-world sites, we construct a more diverse test set than Mind2Web and WebLINX using 500 held-out demonstrations produced by our pipeline. Shown in Figure 11, we train agents using all human data in the training sets for WebLINX and Mind2Web, and compare the performance with agents trained on 80% human data, and 20% data from our pipeline. Agents trained with our data achieve comparable performance to agents trained purely on human data on the official test sets for the WebLINX and Mind2Web benchmarks, suggesting that when enough human data are available, synthetic data may not be necessary. However, when evaluated in a more diverse test set that includes 500 sites not considered by existing benchmarks, agents trained purely on existing human data struggle to generalize. Training with our data improves generalization to these sites by +149.0% for WebLINX agents, and +156.3% for Mind2Web agents, with the largest gains in generalization Step Accuracy appearing for harder tasks.
7 Conclusion
We present a pipeline that unlocks new scaling potential for language model agents by generating internet-scale web navigation data for diverse sites. We employ pretrained language models to generate, attempt, and evaluate web navigation tasks at scales beyond what humans can manually annotate. Our results show that language models outperform human annotators at detecting safe websites, and proposing feasible web navigation tasks. In addition, we find that language models can judge successful web navigation trajectories with an accuracy up to 93.1% for their most confident predictions. We scale task generation to the top 1M sites in the CommonCrawl PageRank, and test agents on 150k live web navigation tasks. Agents based on Llama 3.1 70B solve 16.7% of these tasks zero-shot. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data. When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
Our work reveals several exciting directions for future work. First, our work can be scaled further. The latest CommonCrawl release contains data for more than 300 million sites, suggesting another 1,000 times more data than we explored could be available for training agents by scaling the pipeline. In addition, our language model judge was employed offline, and its high accuracy suggests that it could be used to guide an online algorithm. Finally, we considered primarily text-based agents in this work, and our pipeline could be extended to generate data for multimodal tasks.
8 Acknowledgments
The authors recognize funding via an unrestricted gift from Amazon. In addition, the leading author was partially funded by the US Department of Defense via the NDSEG Fellowship. We thank fellow Ph.D. students who provided feedback on versions of this manuscript, including Michael Luo, Justin Wong, Tiffany Min, Murtaza Dalal, Jing-Yu Koh, Boyuan Chen, and others. We also thank faculty who provided feedback and direction for parts of this work, including Aviral Kumar, Graham Neubig, Shuyan Zhou, and others.
Appendix A Limitations & Safeguards
Language model agents present unique challenges and risks when applied to live tasks on the internet. For instance, agents visiting shopping sites can influence the statistics produced by analytics tools, which can impact prices on products, and product decisions from companies. Furthermore, agents visiting harmful content can add such harmful content to datasets, and perpetuate harmful behaviors into the training data for future agents. We mitigate these risks by carefully designing the task proposal stage of the InSTA pipeline. We consider the risks posed to analytics tools by limiting the engagement between agents and sites. We generate only one task per website, and we limit agents to just 10 actions per site, which includes clicks, typing, dropdown selection actions, and more. By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $30$ seconds of interaction per site on average). By utilizing data from the InSTA pipeline in an offline fashion, as in Section 6 of the main paper, no additional web traffic is generated when training agents. To ensure that agents do not modify the state of the web (i.e. avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide an instruction in the system prompt of the task proposer (see Figure 2) to avoid writing tasks that require the agent to modify the state of the web.
The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate (such as social media, and forums). We explore the performance of the task proposer at filtering out unsuitable sites in Section 4.2, and find that all models detect unsuitable sites with a recall from $0.98$ to $1.0$, and accuracy up to 97%, suggesting our filter is reliable. Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and include harmful, and mature content.
Appendix B Ethical Considerations
One important ethical consideration when gathering data from the internet is to handle copyrighted, private, and sensitive materials carefully. The internet contains vast amounts of personal data created by users that includes personally-identifying-information that should not be included in public datasets. We address this ethical consideration in two ways. First, the task proposer is instructed to filter out social media sites and forums that are likely to contains personally-identifying-information. Second, we store and release only the prompts we used, and traces for agents’ actions—importantly, we do not release any web source code that could be used to recover sensitive data. These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are included in our data, and methods for detecting, replacing, and removing such materials from datasets remains an important task for researchers working on safety.
Appendix C Broader Impacts
As their capabilities broaden, language model agents are being increasingly used to operate real-world systems and APIs. This shift comes with several benefits and risks. Agents that operate your computer to aid in work tasks can significantly boost productivity for certain workers, but can displace others whose jobs have been fully automated. Agents that operate web browsers to complete personal tasks for users can provide convenience, but expose a new attack vector where compromised agents perform unintended actions. Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential in nature, and harder to address purely from a research perspective.
Our data pipeline aims to facilitate internet-scale training for agents using offline data. Training from offline data is important because current agent capabilities are low, and training them on live websites risks taking bandwidth away from real users, while polluting the internet with noisy LLM outputs. Due to these concerns, we advocate for researchers building internet-scale agents to follow the guidelines set forth in this work: (1) agents should not modify the state of the internet by making accounts, creating posts, or submitting content; (2) agents should be throttled to a maximum rate of interaction, and a maximum number of interactions per website; (3) copyrighted, private, and sensitive materials should be removed from training data.
Appendix D Agents.txt & Standards For Internet Agents
Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites—if at all. Desireable controls include rate limits for interactions, limits for maximum numbers of interactions, restrictions to allow agents to interact with certain pages and not others, and restrictions on the kind of data on webpages that agents are allowed to observe (achieved via tagging elements to hide their content from agents). In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from their primary user-facing platform.
Appendix E More Details On Task Generation
You are helping us create tasks for a web navigation system. We will tell you the domain of a website. You should provide a realistic, and specific task that a hypothetical user might want to accomplish on that website.
## Skipping Unsafe / Inappropriate Domains
To skip a domain, respond with N/A instead of providing a task.
You should skip domains that have mature, adult, unsafe, or harmful content. If you are unsure whether a domain is safe, please skip it. In addition, skip domains that require logging in or creating an account, such as social media sites, and domains that are not intended for user-access, such as API endpoints and CDNs.
## Here are some domains to provide tasks for:
* www.amazon.com‘: Find the price of the 24in LG Ultragear Monitor.‘
* www.wikipedia.org‘: Look up the history of the Eiffel Tower on Wikipedia.‘
## Here are some domains to skip:
* fbcdn.net‘: N/A
* api.github.com‘: N/A
Tasks should not require external knowledge, not modify the state of the web, and should not require logging in or creating an account. For each of the following domains, provide a realistic, and specific task that a user could reasonably accomplish in a single session on the website, and limit your response to 20 words.
Figure 12: System prompt for task generation. We carefully design the system prompt for task generation to ensure that sites not suitable for inclusion in the training data for agents are detected and removed. This prompt ensures that proposed tasks are passive in nature, and only involve retrieving information—active tasks like making posts and creating accounts are explicitly not allowed.
We provide the system prompt used for task generation in Figure 12. This prompt was provided to Llama 3.1 70B, GPT-4o, and Gemini 1.5 Pro to generate tasks and filter sites unsuitable for annotation in Section 4. We carefully designed this system prompt to enforce that generated tasks are passive in nature, and do not modify the state of content on the internet. In addition to this system prompt, we employed a list of 100 hand-picked in-context examples of website URLs and appropriate tasks, which are provided in the following JSON list. When querying an LLM to generate tasks, we randomly sample 16 in-context examples from the following list, and provide only these examples to the LLM. This helps promote diverse tasks.
"domain": "archive.org",
"task": "Identify the oldest book available in the public domain on this site."
"domain": "arxiv.org",
"task": "Retrieve the latest preprint paper on machine learning."
"domain": "wikibooks.org",
"task": "Find a freely available textbook on linear algebra."
"domain": "wiktionary.org",
"task": "Get the definition and etymology of the word serendipity’."
"domain": "openlibrary.org",
"task": "Locate an ebook about classic literature that is available for borrowing."
"domain": "openculture.com",
"task": "Find a free online course on ancient history."
"domain": "theguardian.com",
"task": "Retrieve an article discussing recent trends in renewable energy."
"domain": "medium.com",
"task": "Identify a highly rated blog post on productivity hacks."
"domain": "goodreads.com",
"task": "Find the most popular book related to neuroscience."
"domain": "wired.com",
"task": "Retrieve an article about the latest advancements in wearable technology."
"domain": "data.gov",
"task": "Identify the latest government dataset on climate change."
"domain": "kaggle.com",
"task": "Find a well-documented data science competition on image recognition."
"domain": "gov.uk",
"task": "Locate the latest UK government report on healthcare."
"domain": "unsplash.com",
"task": "Find a high-resolution image of the Milky Way Galaxy."
"domain": "pexels.com",
"task": "Retrieve a popular photo tagged with nature’."
"domain": "creativecommons.org",
"task": "Find an article explaining Creative Commons licensing types."
"domain": "pypi.org",
"task": "Retrieve the most downloaded Python package for data analysis."
"domain": "huggingface.co",
"task": "Identify a popular machine learning model on this platform."
"domain": "sciencenews.org",
"task": "Find the most recent article on the health impacts of air pollution."
"domain": "mit.edu",
"task": "Retrieve a publicly available research paper on quantum computing."
"domain": "springer.com",
"task": "Identify the latest edition of a Springer book on robotics."
"domain": "jstor.org",
"task": "Find a research paper discussing the history of the Internet."
"domain": "biorxiv.org",
"task": "Retrieve the most recent bioRxiv preprint on CRISPR technology."
"domain": "medrxiv.org",
"task": "Find a public health preprint related to COVID-19."
"domain": "commons.wikimedia.org",
"task": "Retrieve a high-resolution image of the Eiffel Tower."
"domain": "scholar.google.com",
"task": "Find the most cited article by a specific researcher."
"domain": "plos.org",
"task": "Locate the latest research paper on gene editing published here."
"domain": "flickr.com",
"task": "Find a photo that has been released under a Creative Commons license."
"domain": "datacite.org",
"task": "Retrieve metadata for a dataset related to environmental studies."
"domain": "orcid.org",
"task": "Find the ORCID ID of a well-known researcher in AI."
"domain": "zotero.org",
"task": "Retrieve an article discussing citation management tools."
"domain": "github.com",
"task": "Find the most starred repository on deep learning."
"domain": "figshare.com",
"task": "Retrieve an open dataset on climate patterns."
"domain": "zenodo.org",
"task": "Find the latest publication on open science practices."
"domain": "worldcat.org",
"task": "Locate a catalog entry for a rare book on botany."
"domain": "biodiversitylibrary.org",
"task": "Retrieve a scanned copy of an 18th-century botanical illustration."
"domain": "genome.gov",
"task": "Find the latest update on the Human Genome Project."
"domain": "merriam-webster.com",
"task": "Retrieve the definition and usage of the word quantum’."
"domain": "stanford.edu",
"task": "Find the most recent online lecture on artificial intelligence."
"domain": "edx.org",
"task": "Retrieve a TED Talk on leadership in technology."
"domain": "ted.com",
"task": "Find the latest ocean temperature data available."
"domain": "noaa.gov",
"task": "Retrieve a dataset related to consumer behavior."
"domain": "data.world",
"task": "Find a course on data visualization."
"domain": "curious.com",
"task": "Retrieve a well-cited article on the psychological impact of social media."
"domain": "theconversation.com",
"task": "Identify a recent research paper on biodiversity conservation."
"domain": "nature.com",
"task": "Retrieve the latest article on genomics research."
"domain": "pnas.org",
"task": "Find a science news article on robotics advancements."
"domain": "sciencedaily.com",
"task": "Identify the top story on global health issues."
"domain": "bbc.com",
"task": "Retrieve a recent podcast episode about space exploration."
"domain": "npr.org",
"task": "Locate the most recent update on the global biodiversity status."
E.1 Details For Safety Experiment
This list of examples is also provided in our code release, alongside the script that we used to generate task proposals for the top 1M sites in the CommonCrawl PageRank (The Common Crawl Foundation, 2024). Using these prompts for task generation, we can filter our sites that are unsuitable for annotation, due to containing harmful content, or sensitive user data. To evaluate the performance of our filter, we employed a set of 100 curated websites, where 50 are manually verified as safe, and 50 are manually verified as unsafe based on the filtering conditions. These sites were chosen to span popular sites that typical annotators are likely familiar with, and less popular sites that annotators may not already be familiar with.
safe_sites_list = [’dhss.mo.gov’, dizionari.corriere.it’, southgippsland.vic.gov.au’, ds.iris.edu’, lobbycontrol.de’, ’4rsmokehouse.com’, barnsleyfc.co.uk’, wiwi.uni-wuerzburg.de’, uplandca.gov’, lsus.edu’, wpcode.com’, webopedia.internet.com’, tamko.com’, premierchristian.news’, genome.jgi.doe.gov’, burgerking.ca’, thehugoawards.org’, radio.fm’, thevinyldistrict.com’, unilang.org’, raywhitegroup.com’, grapevinetexas.gov’, sanfrancisco.cbslocal.com’, hyde-design.co.uk’, breastcancerfoundation.org.nz’, ludwigsburg.de’, ignitionrobotics.org’, deliverit.com.au’, kodokan.org’, clickstay.com’, searchdatamanagement.techtarget.com’, oceanario.pt’, wentworthpuzzles.com’, catholicworldreport.com’, quizlet.com’, innovation.nhs.uk’, synonyms.reverso.net’, news.siemens.co.uk’, readability-score.com’, co.modoc.ca.us’, cityofmyrtlebeach.com’, loire.gouv.fr’, lawphil.net’, saem.org’, parmigianoreggiano.it’, engaging-data.com’, itf-tkd.org’, aka.education.gov.uk’, ub.uni-kl.de’, mottchildren.org’]
unsafe_sites_list = [’refer to the code release’]
We provide the list of safe sites in the preceding code block, and in the code release, but we provide the list of unsafe sites in just the code release to keep this manuscript clean of harmful content. Using these lists, we conduct an experiment measuring the accuracy, precision, and recall of our safety filter for detecting sites that are not suitable for training agents. In this experiment, we run the task proposer with the same prompts used to scale annotation the top 1M sites in the CommonCrawl PageRank, and we consider a site to be marked positive for unsafe content if the task proposer generates “N/A” rather than predicting a task. To calculate the human baseline for this experiment, we showed participants the system prompt observed by the task proposer, and asked participants to label sites as safe or unsafe using this per-example prompt:
[…]
Human participants were not allowed to visit the URL shown, and had to determine whether the site is safe for annotation purely from their prior knowledge (the same conditions faced by the task proposer). The 100 sites for the safety experiment were shuffled into a uniformly random order to ensure the order of annotation did not bias the human annotators predictions. One human participant was used to obtain the human baseline result in Table 3.
E.2 Details For Reliability Experiments
Similar to the previous safety experiment, we employed human participants to obtain a human baseline for task feasibility. In particular, we showed human participants the system prompt in Figure 12 for the task proposer, and had them write a task for each of the following websites without visiting the URL (the same conditions faced by the task proposer). The following 100 sites were shuffled into a uniformly random order to ensure the participants were not influenced by the order in which sites were shown. After tasks were proposed by participants, and by LLMs, we evaluated the expert feasibility of tasks by manually attempting to complete the tasks proposed by each set of participants, and marking tasks as feasible, or not feasible based on our own ability to complete them. In total, we annotated 400 tasks, which required 8 hours of annotation. One human participant was used to obtain the human baseline result in Table 4.
reliability_sites_list = [’godaddy.com’, chrome.google.com’, apple.com’, support.cloudflare.com’, support.apple.com’, edition.cnn.com’, go.microsoft.com’, google.de’, w3.org’, yandex.ru’, bfdi.bund.de’, microsoft.com’, apps.apple.com’, networksolutions.com’, support.mozilla.org’, yelp.com’, cnn.com’, ec.europa.eu’, developer.mozilla.org’, icann.org’, books.google.com’, globenewswire.com’, onlinelibrary.wiley.com’, gnu.org’, slideshare.net’, metacpan.org’, porkbun.com’, oag.ca.gov’, spiegel.de’, linuxfoundation.org’, help.opera.com’, mayoclinic.org’, podcasts.apple.com’, nhs.uk’, addons.mozilla.org’, google.fr’, pewresearch.org’, finance.yahoo.com’, weforum.org’, g2.com’, savethechildren.org’, news.com.au’, biblia.com’, yr.no’, engadget.com’, microsoftstore.com’, ema.europa.eu’, theintercept.com’, princeton.edu’, foodandwine.com’, sfgate.com’, voguebusiness.com’, ourworldindata.org’, livingwage.org.uk’, cms.law’, msdmanuals.com’, websitesetup.org’, support.xbox.com’, treehugger.com’, tripadvisor.com.pe’, mondragon.edu’, greenparty.ca’, aaojournal.org’, restaurantpassion.com’, iwillteachyoutoberich.com’, moneyconvert.net’, gesundheitsinformation.de’, ovc.uoguelph.ca’, zdnet.be’, oxfordamerican.org’, snackandbakery.com’, journals.uic.edu’, confused.com’, standards.globalspec.com’, onlyinyourstate.com’, ahsgardening.org’, wyze.com’, nornickel.ru’, viessmann.fr’, benetton.com’, firecomm.gov.mb.ca’, executedtoday.com’, eukn.eu’, fraeylemaborg.nl’, verizon.com/about/news-center’, orthodoxalbania.org’, cheapjoes.com’, bake-eat-repeat.com’, plattformpatientensicherheit.at’, hifinews.com’, cellsignal.com’, thenotariessociety.org.uk’, chosenfoods.com’, westerndressageassociation.org’, pridesource.com’, northtacomapediatricdental.com’, strade-bianche.it’, pvdairport.com’, institute.sandiegozoo.org’, raintaxi.com’]
[…]
E.3 Automatic Task Categorization
You are a helpful scientific assistant categorizing tasks on the web. You will observe a domain and web navigation task, and you should provide a concise categorization of the task in 3 words or less. For example, if the domain is "google.com" and the task is "find a recipe for mashed potato", you may categorize the task as "recipe search".
## Task Format
Here is the format for the task:
[domain]: [task]
Here is what each part means:
‘[domain]‘: The domain of the website you are observing.
‘[task]‘: The task a user is trying to accomplish on the website.
## Response Format
Respond with a category name for the task in 3 words or less, and provide only the category name, do not provide an explanation or justification for the categorization.
Here is the next task, please follow the instructions carefully.
Figure 13: System prompt for task categorization. We employ Llama 3.1 70B to automatically label task categories for our dataset. We prompt the LLM to assign categories in 3 words or less, and set the sampling temperature to $0.5$ to encourage predictions to use more consistent language. Using these categories, we seek to understand agent performance by category.
To better understand the statistics of generated tasks, we employ Llama 3.1 70B to assign task categories. We prompt Llama 3.1 70B with the system prompt in Figure 13 to assign a category in 3 words or less to encourage simple categories. Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean. There is occasional overlap between categories, which can be observed in Figure 14, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated. We provide our task categorization script in the official code release.
Appendix F Understanding Agent Capabilities & Limitations
Figure 14: Largest categories for internet-scale task generation. We assign categories to 150k web navigation tasks generated by our pipeline in Section 4, and visualize the number of tasks for each of the largest 70 categories. Top categories include article search, news search, recipe search, product lookup, and more. The top 12 task categories have more than 1600 tasks assigned to each of them, the mean number of tasks per category is 16.9, and 89% of categories (7741 in total) have fewer than the mean number of tasks.
Figure 15: Most successful categories for internet-scale task generation. We explore the rates of task completion for the top categories of tasks generated by our pipeline. We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 of such categories. Results show that 22 categories are solved with more than a 50% rate with agents based on Llama 3.1 70B.
To complete the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently. Shown in Figure 15, we plot the average judge success probability prediction $\mathbf{r}_{T}$ versus task category for the top 70 most successful categories that have at least 100 tasks assigned to them. Based on the figure, top categories include search for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research. Based on these results, the top 22 categories are solved with more than a 50% rate using agents based on Llama 3.1 70B running zero-shot. As stronger models are developed, the success rates for agents running in our pipeline are likely to improve, and the quality of the data we generate will jointly improve.
In addition to studying the best-performing categories, we also explore the limitations of current agents via their least successful categories. Shown in Figure 16, we select the bottom 70 categories in terms of their average judged success probability for categories with at least 100 tasks assigned to them. Many of these categories require agents to remember and reason about previous interactions, such as the product comparison category. For this category, an agent must review several products, and compare their details from memory. In these cases, access to a note-taking app may improve performance. Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening and playing files. While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment could unlock these abilities in future work.
Figure 16: Least successful categories for internet-scale task generation. Similar to the previous figure, we explore the rates of task completion for the bottom 70 categories that have at least 100 tasks assigned to them. While the majority of the least successful categories have success rates greater than 20%, performance drops as low as 5%. Many of the categories shown in the plot above involve actions that are not feasible given the current limitations of the Playwright API, and may be possible in future work that extends agents to a fully-operable virtual computer environment. In addition, better LLM backbones are likely to improve performance.
Appendix G Agent & Judge System Prompts
We provide the system prompt used with our agent below. This prompt is released in our official code, alongside the observation processor that maps webpage DOM to a compact markdown format, referenced in the system prompt.
You are a helpful assistant operating my web browser. I will show you webpages formatted in markdown, and I want your help to complete a web navigation task. Read the webpage, and respond with an action in JSON to interact with the page, and help me complete the task.
## Formatting The Response
Respond with actions in the following JSON schema:
‘‘‘json
"action_key": str,
"action_kwargs": dict,
"target_element_id": int
Here is what each key means:
- action_key‘: The action to perform.
- action_kwargs‘: Named arguments for the action.
- target_element_id‘: The id of the element to perform the action on.
## Available Actions
Im using playwright, a browser automation library, to interact with the page. Im parsing the value assigned to action_key into a method call on the page object, or an element object specified by the value assigned to target_element_id‘. Here are the available actions:
### Click Action Definition
- click‘: Click on an element specified by target_element_id‘.
### Example Click Action
Suppose you want to click the link ‘[id: 5] Sales link‘:
‘‘‘json
"action_key": "click",
"action_kwargs": {},
"target_element_id": 5
### Hover Action Definition
- hover‘: Hover over an element specified by target_element_id
### Example Hover Action
Suppose you want to hover over the image ‘[id: 2] Company Logo image‘:
‘‘‘json
"action_key": "hover",
"action_kwargs": {},
"target_element_id": 2
### Fill Action Definition
- fill‘: Fill an input element specified by target_element_id with text.
- value‘: The text value to fill into the element.
### Example Fill Action
Suppose you want to fill the input ‘[id: 13] "Name..." (Enter your name text field)‘ with the text John Doe‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "John Doe"
"target_element_id": 13
### Select Action Definition
- select‘: Select from a dropdown element specified by target_element_id‘.
- label‘: The option name to select in the element.
### Example Select Action
Suppose you want to select the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
‘‘‘json
"action_key": "select_option",
"action_kwargs": {
"label": "red"
"target_element_id": 67
### Go Back Action Definition
- go_back‘: Go back to the previous page (‘target_element_id must be null).
### Example Go Back Action
‘‘‘json
"action_key": "go_back",
"action_kwargs": {},
"target_element_id": null
### Goto Action Definition
- goto‘: Navigate to a new page (‘target_element_id must be null).
- url‘: The URL of the page to navigate to.
### Example Goto Action
Suppose you want to open google search:
‘‘‘json
"action_key": "goto",
"action_kwargs": {
"url": "https://www.google.com"
"target_element_id": null
### Stop Action Definition
- stop‘: Stop the browser when the task is complete, or the answer is known.
- answer‘: Optional answer if I requested one.
### Example Stop Action
‘‘‘json
"action_key": "stop",
"action_kwargs": {
"answer": "Im done!"
"target_element_id": null
Thanks for helping me perform tasks on the web, please follow the instructions carefully. Start your response with an explanation in 50 words, and choose exactly one action you would like to perform.
We also provide the system prompt used with out LLM judge. The system prompt instructs the judge to predict a json-formatted dictionary that contains a “success” key, and an “on_right_track” that represent the estimated probability that the task is successful, and that the agent is on the right track towards solving the task, respectively. These distinctions are adapted from Koh et al. (2024b), and help us filter for high-quality training data by distinguishing trajectories that were solved by the agent’s own actions from trajectories that were solved by chance.
You are a helpful assistant providing feedback on a web automation script. I will show you a list of previous actions, the current webpage formatted in markdown, and the proposed next action. I want your help evaluating the proposed action, to determine if the desired task is complete, or if we are on the right track towards future completion.
## Reading The Action Schema
You will see actions in the following JSON schema:
‘‘‘json
"action_key": str,
"action_kwargs": dict,
"target_element_id": int
Here is what each key means:
- action_key‘: The action to perform.
- action_kwargs‘: Dictionary of arguments for action.
- target_element_id‘: The id of the element to perform the action on.
## Available Actions
Im using playwright, a browser automation library, to interact with the page. Im parsing the value assigned to action_key into a method call on the page object, or an element specified by the value assigned to target_element_id‘. Here is an example action:
### Example Click Action
Here is an example where the script clicked the link ‘[id: 5] Sales link‘:
‘‘‘json
"action_key": "click",
"action_kwargs": {},
"target_element_id": 5
### Example Select Action
Here is an example where the script selected the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
‘‘‘json
"action_key": "select_option",
"action_kwargs": {
"label": "red"
"target_element_id": 67
### Example Goto Action
Here is an example where the script opened google search:
‘‘‘json
"action_key": "goto",
"action_kwargs": {
"url": "https://www.google.com"
"target_element_id": null
### Example Stop Action
Here is an example where the script stopped with the message "Im done!":
‘‘‘json
"action_key": "stop",
"action_kwargs": {
"answer": "Im done!"
"target_element_id": null
## Formatting The Response
Think step by step, and start your response with an explanation of your reasoning in 50 words. Then, provide an evaluation in the following JSON schema:
‘‘‘json
"success": float,
"on_right_track": float,
Here is what each key means:
- success‘: What is the probability the desired task has been completed successfully, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
- on_right_track‘: What is the probability the script is on the right track towards a future success, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
Thanks for helping me evaluate the script, please follow the instructions carefully. Start your response with a step by step explanation. Then, provide an evaluation in the JSON schema above.
Appendix H Details For Training Agents
To understanding the utility of the generated data for training agents, we filter the data, and compare our filtered data to human demonstrations on the WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023) benchmarks. In particular, we sweep over different sizes of random subsets of human actions, from 32 to 256, which helps us understand the value of synthetic data generated from the InSTA pipeline versus different scales of human data. We then fine-tune models based on google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks. We employ identical training hyperparameters to those used in Lù et al. (2024) and Deng et al. (2023) to ensure that our results are comparable to previous work. Section H reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites.
In order to prepare our data, we employ three filtering rules. In the first rule, we filter for data where the agent was predicted to have succeeded at the task with conf = 1, and was predicted to be on the right track with conf = 1. This filtering rule is motivated by our findings in Section 5, where we found that our LLM judge based on Llama 3.1 70B has an accuracy up to 93.1% at detecting successful trajectories for its predictions with conf = 1. Filtering based on both “success” and “on_right_track” conditions is essential to obtain data where the agent directly caused the task to be solved, rather than the task being solved by external conditions. The next filtering rule we use is to select trajectories with at least three actions, which helps create training data that is not too easy (i.e. not solved after just one or two actions). Finally, we select tasks where the agent did not encounter any errors during execution. These include being presented with server errors such as 404 Not Found, and 403 Forbidden, encountering a captcha, and being blocked, even if just temporary, from the target website. These filtering steps produce an automatically curated set of $7,463$ demonstrations from our pipeline where agents successfully completed tasks generated by the InSTA pipeline. We reserve 500 demonstrations from this pool for our test set, and the rest for training agents in Figure 10. The original Mind2Web dataset contains $2,350$ tasks.
Appendix I Additional Related Works
While writing this paper, concurrent work was released that introduces a Proposer-Agent-Evaluator framework for web navigation agents (Zhou et al., 2024c). There are several key differences between our work and theirs, and the most important difference is scale. We generate tasks for 1M sites on the internet, whereas their work considers just 85 real-world sites, 5 sites from WebArena (Zhou et al., 2024b), and 13 sites from WebVoyager (He et al., 2024). The second difference is evaluation. Safety and reliability play crucial roles when gathering data, and we conduct an analysis on the safety and reliability of data generated by our method on 100 real-world sites. Another major difference pertains to offline learning. Offline learning should be used when scaling agents because current agent capabilities are low, and training them online risks polluting the internet with noisy LLM outputs, while taking bandwidth away from real users. The final difference pertains to the train-test split. We train agents on diverse internet data, and transfer to target benchmarks, while the agents presented in Zhou et al. (2024c) train on sites from target benchmarks using synthetic tasks. Our train-test split is stronger, and evaluates the ability for agents trained on our synthetic data to generalize to novel websites, domains, and tasks.
Appendix J Hyperparameters
We provide a list of the hyperparameters used in this work in Table 1. Values are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to employ standard hyperparameters for training agents on WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023). On the Mind2Web benchmark, we did not observe a loss in performance for training longer on few-shot experiments, and kept training iterations fixed between few-shot and all-available-data experiments. This is likely due to the FLAN-T5 model used in Mind2Web being too small to overfit to the data. For experiments on WebLINX, we found that performance for models trained on just human data degrades if models are trained for the full $10,000$ steps, so a smaller $2,000$ steps was chosen.
Hyperparameter Name Value
OpenAI Model Name gpt-4o
Google Model Name gemini-1.5-pro
Llama Model Name meta-llama/Llama-3.1-70B-Instruct
CommonCrawl PageRank Revision cc-main-2024-apr-may-jun-host-ranks.txt.gz
Number of sites before filtering $1,000,000$
Number of tasks after filtering $146,746$
Max Tokens Per Observation $4,096$
Max Tokens Per Action $2,048$
Max Tokens Per Judgement $2,048$
Max Tokens Per Task $64$
Max Observations Per Agent Context 5
Max Actions Per Agent Context 5
Max Observations Per Judge Context 1
Max Actions Per Judge Context 5
OpenAI Inference API Sampling Temperature 0.5
OpenAI Inference API Sampling Top P 1.0
Mind2Web HuggingFace Model Name google/flan-t5-large
Mind2Web Few-Shot Training Iterations $11,505$
Mind2Web Training Iterations $11,505$
Mind2Web Batch Size $32$
Mind2Web Learning Rate 5e-5
WebLINX HuggingFace Model Name meta-llama/Llama-3.1-8B-Instruct
WebLINX Few-Shot Training Iterations $2,000$
WebLINX All Training Iterations $10,000$
WebLINX Batch Size $16$
WebLINX Learning Rate 5e-5
Filtering Success Threshold 1.0
Filtering On Right Track Threshold 1.0
Few-Shot $p_{\text{real}}$ 50%
All Human Data $p_{\text{real}}$ 80%
Table 1: Hyperparameters used in our paper. We organize hyperparameters into five sections, including names of language model backbones, parameters of the data generation pipeline, sampling parameters for the OpenAI inference API, training parameters used by corresponding benchmarks, and filtering parameters used to prepare our data for training agents in Section 6.
Appendix K Cost Analysis For Llama 3.1 70B
To contextualize how Llama 3.1 70B is important for a project at this scale, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models. As the analysis shows, using Llama 3.1 70B is most feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise in accuracy and performance. We have deep gratitude for the Llama team at Meta working to make developments in language modeling available to the research community at no cost.
Variable Name Value
Number of tasks $146,746$
Typical tokens per observation $1,024$
Max observations per agent context window $5$
Typical agent / judge response size $128$
Max tokens per system prompt $1,024$
Max steps per task $10$
Estimated tokens processed by the agent
$146,746*(((1,024+128)*5+1,024)*10-(1,024+128)*15)=$ $7,419,477,760$
Tokens processed by the judge
$146,746*(1,024+1,024+128*10)=$ $488,370,688$
Total tokens processed $7,907,848,448$
Expected API cost for GPT-4o \$ $39,539.24$
Expected API cost for Gemini 1.5 Pro \$ $9,489.42$
Expected AWS compute cost for serving Llama 3.1 70B
       (14 days for two 8-gpu v100 spot instances) \$ $6,622.56$
Percent saved using Llama 3.1 70B $[30.2,83.3]$
Table 2: Cost analysis for different LLM models in the fully-scaled pipeline. This table provides statistics for the number of tokens that were processed by our pipeline, and why serving using a local LLM engine like vLLM is important for bringing down costs.

### v2

@@
-Towards Internet-Scale Training For Agents
+InSTA: Towards Internet-Scale Training For Agents
 Abstract
-The predominant approach for training web navigation agents gathers human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data are an inefficient resource.
+The predominant approach for training web navigation agents is to gather human demonstrations for a set of popular websites and hand-written tasks, but it is becoming clear that human data is an inefficient resource.
 We develop a pipeline to facilitate internet-scale training for agents without laborious human annotations.
-In the first stage, an LLM generates tasks for 150k diverse websites.
+In the first stage, an LLM annotates 150k sites with agentic tasks.
 In the next stage, LLM agents complete tasks and produce trajectories.
-In the final stage, an LLM reviews the trajectories and judges their success.
-Language models are competitive with human annotators, detecting and filtering out harmful content with an accuracy of 97%, generating feasible tasks with an 89% rate, and judging successful trajectories with an 82.6% accuracy.
-Scaling the pipeline, agents based on Llama 3.1 70B solve 16.7% of tasks for 150k sites.
-Training on the data generated by our pipeline is competitive with training on human demonstrations.
-In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data.
-When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
-Code will be available at: data-for-agents.github.io.
+In the final stage, an LLM filters trajectories by judging their success.
+Language models are powerful data curation tools, identifying harmful content with an accuracy of 97%, judging successful trajectories with an accuracy of 82.6%, and producing effective data.
+We train agents based on Qwen 3 1.7B that are competitive with frontier LLMs as web agents, while being smaller and faster.
+Our top agent reaches a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash.
+We are releasing code, models and data at: https://data-for-agents.github.io.
 1 Introduction
-The predominant approach to training LLM-based web navigation agents is to collect human demonstrations on a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023).
-Human data can be laborious to collect and becomes costly to scale as the breadth of skills that users require from language model agents grows.
-There are more than three-hundred million sites on the Internet (The Common Crawl Foundation, 2024), and the range of sites that researchers can manually prepare for human annotation represents a tiny fraction of the Internet.
-The key problem is that human data can become unreliable at scale.
-Human-written web navigation tasks are highly effective for popular sites, but reliability drops for sites with lower popularity due to annotators’ lack of familiarity.
-For these sites with lower popularity, which represent the majority of sites on the internet, human-written web navigation tasks are feasible only 40% of the time (discussed in Section 4), requiring a costly manual verification step.
-For the same pool of sites, language models improve feasibility rates to more than 80%.
-There is a growing need to automate data pipelines for the next generation of agents trained at internet scale.
-We address a key challenge by reducing the dependency on human data in the agent pipeline.
-We develop an automatic data pipeline that aims to facilitate Internet-Scale Training for Agents, which we shorten to InSTA, a pipeline that relies on synthetic web navigation tasks proposed, attempted, and evaluated by language models.
-Figure 1: Overview of the proposed agent pipeline.
-We develop a pipeline for training web navigation agents at internet scale using tasks proposed, attempted, and evaluated by pretrained large language models.
-We generate 150k diverse tasks across 1M internet sites.
-Code for our data generation pipeline, and traces for agent rollouts will be available on our website: data-for-agents.github.io.
-Our method operates in three stages.
-In the first stage, we employ a language model to propose candidate web navigation tasks for an agent to perform across 150k live sites on the Internet.
-Current works are limited to 200 popular sites (Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023) that human annotators are likely to be familiar with.
-Language models help us scale to 1,000 times more sites than current efforts, with better coverage of real-world sites.
-One major consideration when scaling up training for agents is safety: building safe agents requires that we avoid sites with harmful, unsafe, or dangerous content.
-We evaluated the ability of language models to detect such content and aggressively filtered out 85% candidates from the initial 1M sites to 150k sites that are judged safe by language models.
-These models succeed at detecting safe content with an accuracy of 97%, compared to 75% human accuracy.
-With tasks generated across a safe and diverse set of websites, we proceed to run language model agents to attempt the generated tasks.
-In the second stage of the pipeline, a language model agent attempts to complete tasks using a web browser.
-We provide the entire Playwright API to the agent, which operates the browser by generating function calls in the Playwright API, a standard choice in recent web navigation benchmarks (Chezelles et al., 2024; Du et al., 2024; Drouin et al., 2024).
-While existing LLM agents can be directly applied to our tasks, measuring their success presents another challenge: sites organize data in bespoke formats.
-Previous efforts circumvent this by enlisting human labelers to review attempts manually and judge their success (Lù et al., 2024; Deng et al., 2023), or by designing sites for benchmarking where formats are consistent (Zhou et al., 2024b; Koh et al., 2024a).
-However, neither approach scales efficiently, and human annotators may not be reliable for sites they are not familiar with in the tail of the distribution.
-In the third stage of the pipeline, we scale the evaluation using language models.
-We employ LLMs to judge (Lightman et al., 2024) whether a task is solved by the final timestep and obtain an accuracy of up to 93.1% in detecting successful trajectories for the most confident predictions.
-Llama-3.1-70B-Instruct solves 16.7% of zero-shot tasks with a judge confidence of conf = 1.
-In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1%, respectively, when training agents on mixtures of data from our pipeline and human data.
-When training agents with all human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
-Our work develops InSTA, an automated data pipeline that aims to facilitate Internet-scale training for agents.
-Driven by pretrained language models, we generate web navigation tasks for 150k live websites, run LLM agents, and judge their success.
-By removing humans from the agent pipeline, we can scale and extend coverage towards sites where human data is less reliable.
-We present an analysis to improve safety and show that our data can outperform human demonstrations for training agents.
-Code for the pipeline will be available at: data-for-agents.github.io.
+The predominant approach for training LLM web navigation agents is to collect human demonstrations for a set of manually curated websites and tasks (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023).
+Human data can be laborious to collect, and becomes costly to scale as the breadth of skills that users require from language model agents grows.
+There are more than 300M sites on the western internet according to The Common Crawl Foundation (2025), and the range of sites that researchers have annotated represents a tiny fraction of the vast available data.
+And crucially, the existing human data is static.
+There is a growing need to automate pipelines for training the next generation of language model agents in a dynamic internet-scale environment.
+This paper addresses the core challenge of building this environment—reducing dependence on human annotations.
+We develop an automatic pipeline that aims to facilitate internet-scale training for agents, which we refer to as the InSTA pipeline.
+Figure 1: Overview of the InSTA pipeline.
+Our work unlocks a dynamic internet-scale environment that allows training small models to match frontier LLMs as agents, on a fraction of the budget.
+Starting from the top 1M sites on the internet, we annotate 150k sites with challenging agentic tasks, and release the entire pipeline, including code, models and an official huggingface dataset, on our website: https://data-for-agents.github.io.
+The InSTA pipeline has three stages.
+In the first stage, we employ a language model task proposer that annotates 150k sites with live web navigation tasks for agents to perform.
+Existing works are limited to 200 popular websites (He et al., 2024; Lù et al., 2024; Rawles et al., 2023; Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025) that researchers have annotated manually, and a handful of synthetic websites (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a).
+Our first goal in this paper is to improve coverage of real-world sites.
+To accomplish this, we cast a wide net.
+Starting from the top 1M sites on the internet ranked by popularity, our task proposer filters down to 150k sites that have safe content.
+Safety is critical when building autonomous agents, and our task proposer succeeds at detecting harmful content with an accuracy of 97%.
+Tasks are generated for sites marked as safe by the task proposer, and we run language model agents to complete the generated tasks.
+Agent progress is then fed back to the task proposer, which reviews trajectories and judge evaluations in order to assign a harder task based on the latest content of the website, closing the loop.
+Scaling the task generation loop, we annotate 150k diverse sites with challenging agentic tasks, and release an official huggingface dataset: data-for-agents/insta-150k-v2.
+Motivated by the importance of internet data to progress in modern deep learning, our second goal in this paper is an internet-scale data flywheel for training LLM agents.
+We approach this by harnessing LLMs as data curation tools.
+After generating tasks, the pipeline employs pretrained LLMs as agents to complete tasks and produce trajectories, which are filtered by a judge to select the best data.
+Agents control a virtual web browser and produce reasoning traces that contain function calls to interact with and navigate live webpages.
+The judge produces a reasoning trace that considers whether a trajectory is successful, and scores the agent on a continuous scale from 0 to 1.
+We scale the pipeline to create a large reasoning dataset for multimodal agents, with 2.2M screenshots, 2.2M traces for actions, and 150k traces for the judge.
+Our data unlocks great potential in small language models as agents.
+We train a series of models based on Qwen 3 1.7B on varying scales of data from the InSTA pipeline, and match the performance of frontier LLM agents, on a fraction of the budget.
+Our top agent has a success rate of 56.9%, outperforming the data collection policy Qwen 3 235B, beating a 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash.
+To share our progress, we are releasing the entire pipeline, including code, models and data, on our website: https://data-for-agents.github.io.
 2 Related Works
 Language Model Agents.
-There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b) as the backbone for agents (Andreas, 2022).
-These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where text formats differ significantly from their training data.
+There is an emerging paradigm in modern NLP using language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b) as the backbone for agents (Andreas, 2022).
+These models show impressive reasoning capabilities (Bubeck et al., 2023; Zhong et al., 2024; Valmeekam et al., 2024) that allow them to generalize to downstream applications, such as web navigation, where observations differ from LLM training data.
 Search algorithms provide a secondary axis to improve the reasoning capabilities of language model agents (Yao et al., 2023b; Besta et al., 2024; Koh et al., 2024b; Zhou et al., 2024a) by providing an explicit algorithmic scaffold and allowing test-time compute to improve reasoning steps (Snell et al., 2024; Zhong et al., 2024).
-Although most of the work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024) as target benchmarks are becoming more difficult for zero-shot language models.
+Although most recent work focuses on running language models as zero-shot agents, fine-tuning language models to improve their effectiveness as agents is becoming popular (Putta et al., 2024; Zeng et al., 2023; Zhang et al., 2023; Hong et al., 2023; Xie et al., 2024; Wang et al., 2024; Zhou et al., 2024c; Murty et al., 2025) as target benchmarks are becoming more difficult to solve zero-shot.
 Agent Pipelines.
-There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024).
-However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023), or the same tasks (Putta et al., 2024).
-We instead consider a setting where tasks and environment configurations are entirely separate between training and testing, creating a strong train-test split that follows recommended practice.
-This presents a challenge: the web navigation data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024).
-We address this challenge by increasing the scale of data collection and better coverage of the distribution of real-world sites.
-We train on diverse tasks generated by our pipeline and successfully transfer agents trained on our data to downstream benchmarks while maintaining a strong train-test split.
-Our training procedure resembles a modified FireAct (Chen et al., 2023), where language models jointly propose and evaluate tasks for agents.
+There are a growing number of agent pipelines aimed at fine-tuning language models to improve their effectiveness as agents (Mitra et al., 2024; Zeng et al., 2023; Putta et al., 2024; Chen et al., 2023; Ou et al., 2024; Zhou et al., 2024c; Murty et al., 2025).
+However, driven by the limited data available, many such works train on data with significant overlap with their test environment, either with different tasks for the same environment configuration as the test setting (Deng et al., 2023; Zhou et al., 2024c; Murty et al., 2025), or even the same tasks (Putta et al., 2024).
+We consider a setting where tasks and environment configurations (i.e.
+websites) are entirely separate between training and testing, creating a strong train-test split.
+This presents a challenge: human data for training LLM agents is limited (Deng et al., 2023; Lù et al., 2024).
+We address this challenge by reducing dependence on human data in agent pipelines.
+We employ LLMs as data curation tools to automatically design challenging tasks, and select the best training data.
+Our pipeline allows us to train small models that match frontier LLMs on Web Voyager (He et al., 2024), without using any data from Web Voyager.
+Contrast this with methods that train primarily on data from Web Voyager (Zhou et al., 2024c; Murty et al., 2025), and may not transfer to other benchmarks (Xue et al., 2025).
 Agent Datasets.
-The majority of datasets for training web navigation agents rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024).
-This approach has limits, as the breadth and diversity of tasks researchers can manually curate are dwarfed by the volume of sites on the Internet.
-There are more than three-hundred million sites on the Internet according to The Common Crawl Foundation (2024), and existing datasets are limited to about 150 popular sites that human annotators are already familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024).
-There are hypothetical 1,000,000 times more data that could be available if we can efficiently harness this previously untapped resource.
-However, most sites are relatively obscure and human annotators are unreliable for sites they are not already familiar with.
-Finding suitable annotators becomes impractical at this massive scale, so we adapt language models to propose, attempt, and evaluate web navigation tasks.
-Although we are not the first to consider synthetic data for training agents (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), we have developed a key approach to harness internet-scale data efficiently.
+Datasets for training web navigation agents typically rely on human annotators to create tasks (Zhou et al., 2024b; Koh et al., 2024a; Rawles et al., 2023), and provide demonstrations for tasks (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Shen et al., 2024).
+However, the amount of data researchers have annotated represents a tiny fraction of the available internet data.
+There are more than 300M sites on the internet according to The Common Crawl Foundation (2025), yet existing datasets are limited to 200 popular sites that human annotators are familiar with (Deng et al., 2023; Lù et al., 2024; Shen et al., 2024).
+Human data can be laborious to collect, and becomes costly to scale as the capabilities users require from agents grows.
+And crucially, human data is static.
+Our work moves away from fixed datasets for training agents, and towards a dynamic internet-scale environment that grows with an ever-changing internet.
+We are not the first to build an environment (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; He et al., 2024), nor are we the first to consider synthetic data (Gandhi et al., 2024; Ou et al., 2024; Setlur et al., 2024; Tajwar et al., 2024), but we have solved a key challenge that unlocks the internet as the largest environment for agents.
 Language Model Judges.
-Core to our pipeline is a language model evaluator.
-Using language models to judge the correctness of responses is becoming popular to improve accuracy for LLMs (Li et al., 2024), and applications include verifying reasoning steps (Zhang et al., 2024), rejection sampling (Snell et al., 2024; Sun et al., 2024), prioritizing frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filtering out harmful responses (Inan et al., 2023), providing feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and providing ratings for alignment (Lee et al., 2024; Ouyang et al., 2024).
-Our use of language models to evaluate agent tasks is inspired by the generative verifier in Zhang et al.
-(2024), and modified from the multimodal verifier in He et al.
-(2024), where our language model predicts a confidence score that a task is solved, which is used to identify successful attempts.
+Using LLMs to judge the correctness of responses is becoming popular to refine LLM predictions (Li et al., 2024), including to verify reasoning steps (Zhang et al., 2024), rejection sample (Snell et al., 2024; Sun et al., 2024), prioritize frontier nodes in search algorithms (Zhou et al., 2024a; Koh et al., 2024b), filter out harmful responses (Inan et al., 2023), provide feedback for response improvement (Madaan et al., 2023; Paul et al., 2024; Patel et al., 2024; Yuksekgonul et al., 2024), and provide ratings for alignment (Lee et al., 2024; Ouyang et al., 2024).
+Our use of language models to evaluate agents is inspired by Generative Verifiers (Zhang et al., 2024), and the multimodal verifier in He et al.
+(2024).
+One difference is our verifier predicts a reasoning trace that scores the agent from 0 to 1, which helps us rank trajectories to select the best data.
 3 Language Model Agents
-Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a web environment in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$.
-Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{sys}}$.
-$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}\big{(}\;\text{LLM}([\;\mathbf{x}_{\text{sys}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;])\;\big{)}$ (1)
+Language model agents are a class of decision-making agents represented by $\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\mathbf{c})$, a policy that processes multimodal observations $\mathbf{s}_{t}$ (from a virtual web browser in our case) and predicts textual actions $\mathbf{a}_{t}$ to complete a task $\mathbf{c}$.
+Underneath this abstraction, a large language model (LLM) generates actions via the next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{agent}}$.
+$\mathbf{a}_{t}=f^{\;\text{text}\to\text{act}}(\;\text{LLM}(\;\mathbf{x}_{\text{agent}},\mathbf{c},\text{Enc}(\mathbf{s}_{t})\;)\;)$ (1)
 Environment representations for observations and actions typically differ from the expected input format of the language model (typically images and text), and functions are introduced that map the observations to a multimodal prompt $\text{Enc}(\cdot)$, and parse actions from the language model generated response $f^{\;\text{text}\to\text{act}}(\cdot)$.
 For web navigation, the environment state $\mathbf{s}_{t}$ is HTML DOM, and is often formatted as raw HTML code, an Accessibility Tree, Set-of-marks, or screenshots (Zhou et al., 2024b; Koh et al., 2024a; Chezelles et al., 2024; Shen et al., 2024).
-In this work, we employ an off-the-shelf HTML-to-Markdown parser to convert observations into a textual format.
+We built a fast Markdown parser that converts webpage observations into a compact readable format (refer to the code).
 Action formats vary between works, and we build on Schick et al.
 (2023)’s function-calling framework, where a language model generates code that is parsed into a function name and corresponding arguments.
-Given a set of strings $L$, and a set of function argument values $G$, the set of actions $\mathcal{A}$ is:
-$\mathcal{A}=\big{(}L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\big{)}$ (2)
-where $L_{\text{func}}$ is the set of function names in our API, and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$).
-We provide agent access to the entire Playwright API (Microsoft, 2024), a Microsoft-developed browser automation library that wraps around a Chromium web browser.
-The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the task is complete, after which point the agent calls stop with an optional answer:
+Given sets of function and argument names $L_{i}$, and sets of argument values $G_{i}$, the action space $\mathcal{A}$ is:
+$\mathcal{A}=L_{\text{func}}\times(L_{\text{arg1}}\times G_{\text{arg1}})\times(L_{\text{arg2}}\times G_{\text{arg2}})\times\cdots\times(L_{\text{argN}}\times G_{\text{argN}})$ (2)
+where $L_{\text{func}}$ is the set of function names on the page object in the Playwright API (Microsoft, 2024), and function arguments have a name and value ($L_{\text{arg1}}\times G_{\text{arg1}}$) corresponding to the Playwright API.
+We allow the agent access to call arbitrary functions in Playwright (Microsoft, 2024), a Microsoft-developed browser automation library that wraps a headless web browser.
+The agent’s goal is to complete a web navigation task specified via a natural language instruction $\mathbf{c}\in L$, starting from an initial URL, and operating the browser via function calls to the Playwright API until the desired task is complete, after which the agent calls the stop function and provides a final response:
 […]
-We prompt the language model backbone to generate actions wrapped in a JSON code block for straightforward parsing.
-The action parser $f^{\;\text{text}\to\text{act}}$ consists of a regex template that matches the first JSON code block, such as the example in Figure 1, followed by JSON decoding of the string contents within the code block.
-When parsing fails due to invalid syntax, we generate a new response until parsing succeeds.
-Equipped with a language model agent that makes calls to the Playwright API, we face a crucial roadblock that impedes scaling: obtaining large and diverse data.
+We prompt the agent to produce a reasoning trace of a desired length (ablated in Figure 10) that contains function call actions as JSON in a fenced code block.
+To parse actions from the response, we employ a regex template that matches the first JSON code block, and a JSON decoder $f^{\;\text{text}\to\text{act}}(\cdot)$ to parse the contents within the code block.
+When parsing fails due to invalid syntax, we allow the agent to generate a second response.
+Equipped with a language model agent that makes function calls with the Playwright API, we may consider the crucial task of obtaining large and diverse data.
 4 Internet-Scale Task Generation
-Building internet-scale agents requires a diverse scaffold of tasks and environment configurations beyond what can be attained via manually curated examples annotated by humans.
-We develop a pipeline to efficiently harness vast quantities of sites on the Internet that aims to facilitate Internet-Scale Training for Agents, shortened to InSTA.
-Our pipeline uses pretrained language models to generate, attempt, and evaluate synthetic web navigation tasks for a more diverse pool of sites than current efforts that rely on tasks designed by researchers (Deng et al., 2023; Zhou et al., 2024b; Putta et al., 2024; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024).
-Human data is a valuable resource, but can be inefficient to gather.
-Our work shows that language models can be as accurate as human annotators.
-By removing human data from the agent pipeline, we can improve the safety and reliability of tasks, and efficiently scale task generation to 1M sites.
+Training the next generation of LLM agents requires a large and diverse set of websites and tasks beyond what researchers have annotated so far (Deng et al., 2023; Zhou et al., 2024b; Koh et al., 2024a; Liu et al., 2024; Lù et al., 2024; Rawles et al., 2023; He et al., 2024).
+We develop an approach to efficiently annotate vast numbers of sites from diverse sections of the internet with agentic tasks.
+Our approach introduces two important desiderata: (1) it should not rely on human annotations, and (2) tasks should derive from a feedback process that deeply explores the environment.
 4.1 Language Model Task Proposer
-In the first stage, we scale the generation of web navigation tasks to 1M diverse websites using a Language Model Task Proposer.
-Shown in Figure 2, the task proposer serves two key functions in the pipeline: (1) filtering out harmful content and websites that cannot be safely annotated, and (2) proposing realistic web navigation tasks that a hypothetical user might want to accomplish.
-Figure 2: Task proposal and filtering for 150k live websites.
+The key idea in stage one of the pipeline is a feedback loop, where a language model task proposer $\psi_{\text{LLM}}(\cdot)$ guides exploration on a website via an initial easy task.
+We then run a language model agent to explore the site, conditioned on the initial task, which produces an exploratory trajectory that is fed back to the task proposer.
+Conditioned on a trajectory that deeply explores the website, the task proposer then creates a harder, grounded task.
+This process is summarized as an equation.
+$\mathbf{c}\sim\psi_{\text{LLM}}(\mathbf{c}|\mathbf{w},\tilde{\mathbf{c}},\underbrace{\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}}_{\text{trajectory}})\;\;\text{st}\;\;\mathbf{a}_{t}\sim\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\tilde{\mathbf{c}})\;\;\mathbf{s}_{t+1}\sim\mathbb{P}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$ (4)
+In this equation, the website url is $\mathbf{w}$, the initial task is $\tilde{\mathbf{c}}$, the trajectory includes the states $\mathbf{s}_{t}$ and the actions $\mathbf{a}_{t}$, and the harder, grounded task is $\mathbf{c}$.
+Highlighted in Figure 2, we annotate 150k sites with tasks by scaling Equation 4, and release them on huggingface at: data-for-agents/insta-150k-v2.
+Figure 2: Annotating 150k live sites with agentic tasks.
 Starting from 1,000,000 websites, we employ a pretrained language model that marks sites as safe/unsafe for annotation, and assigns a realistic task that a hypothetical user might want to accomplish on each site.
@@
 Model Details.
-We utilize pretrained and frozen language models that conform to a chat interface and accept a system prompt $\mathbf{x}_{\text{sys}}$, and a series of in-context examples through interleaved user and assistant prompts $\mathbf{x}_{\text{usr}}$ and $\mathbf{x}_{\text{ast}}$.
-The system prompt used for task generation is shown in Figure 2, and outlines all cases for which sites are considered unsafe for annotation.
-We consider the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b; a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google.
+We utilize pretrained and frozen language models that conform to a chat interface and accept system, user, and assistant prompts.
+The task proposer system prompt is listed in Appendix 11, and details all cases for which sites are considered unsafe.
+We employ the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b, a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google.
 Inference is served using vLLM (Kwon et al., 2023) for the Llama series of models.
-We employ a sampling temperature of 0.5, and a maximum budget of 64 newly generated tokens; all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
+We employ a sampling temperature of 0.5 and a maximum budget of 1024 new generated tokens, while all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
 Prompt Details.
-The goal of the task proposer is to accurately detect unsafe websites and generate realistic web navigation tasks when appropriate.
-We prompt the task proposer with the system prompt in Figure 2, a series of in-context examples (listed in Appendix E), and a final user prompt containing just the URL of the target website.
-We instruct the LLM via the system prompt to provide a task for the target website or to return “N/A” and mark the website as not suitable for annotation.
-This format produces a throughput of 20 websites per second for Llama 3.1 70B served on 16 GPUs with vLLM, processing 1M sites in 14 hours.
-The efficiency of stage one aids in scaling to large numbers of sites on the Internet, but we must not compromise safety and reliability for efficiency.
-To understand the trade-offs presented by our task proposal approach, we compare against typical human annotators at detecting safe websites for annotation and creating realistic agent tasks.
+[…]
+The design of the task proposer as a feedback process is important, and the full potential of this design will be realized in future work that trains agents with an on-policy reinforcement learning algorithm.
+In such future work, the task proposer can be used within the RL loop to generate incrementally harder tasks as agents learn.
+For this paper, we employ one loop of task generation.
+Figure 3: Most frequent words in our tasks.
+This wordcloud shows the top 500 most frequent words in tasks from the training set of our official huggingface dataset.
+The size of each word corresponds to its frequency in the dataset.
+Our tasks span diverse categories and lexicon.
+4.2 Safety & Reliability
+Safety is critical when building autonomous agents.
+The internet contains significant amounts of content that should be removed from training data, in order to avoid agents learning harmful behaviors.
+To understand the robustness of our safety filter, we input 100 carefully selected websites to the task proposer, of which 50 contain harmful, or mature content.
+Table 2 reports the accuracy, precision, and recall of the safety filter on this data.
+For a variety of LLMs, the safety filter displays high accuracy—up to 97% of websites are correctly classified, and recall for detecting unsafe websites is as high as 1.0, suggesting that nearly all unsafe websites are detected by the safety filter.
 Method Acc. Prec. Recall
@@
 Gemini 1.5 Pro 97% 0.96 0.98
-Human Baseline 75% 0.71 0.84
-Figure 3: Accuracy for detecting harmful sites.
-We select 100 website domains, where 50 are safe, and 50 are unsafe based on the criteria in Figure 2.
-Pretrained language models exceed the accuracy and recall of human annotators at detecting harmful sites that are unsuitable for training agents safely.
-4.2 Improving Safety
-Language models beat single pass non-expert annotators at detecting websites suitable for annotation.
-To evaluate detection performance, we employ the task proposer as a classifier and consider sites where the task proposer returns “N/A” as the positive class.
-We curate 50 safe and 50 unsafe domains, based on the filtering conditions outlined in the system prompt in Figure 2 (selected websites and their URLs are listed in Appendix E).
-We generate task proposals for each site and measure the accuracy, precision, and recall of our safety filter compared to human annotators.
-The annotators are asked to classify each site as suitable or unsuitable for annotation based on the website URL, and the criteria listed in the system prompt the same observations given to the task proposer to ensure a fair comparison.
-Results are presented in Table 3.
-Understanding The Results.
-Language models outperform human annotators by 29.3% in accuracy, 35.2% in precision, and 31.0% in recall at detecting harmful sites.
-While larger models like Gemini 1.5 Pro show the best overall accuracy, smaller models like Llama 3.1 70B display high recall with a minor decrease in accuracy.
-Recall matters most for safety filters, and these results suggest that Llama 3.1 70B is sufficient to detect most harmful sites with high confidence.
-4.3 Improving Reliability
-Language models are more reliable than human annotators at creating realistic web navigation tasks.
-To evaluate reliability, we measure the rate at which human workers are able to perform web navigation tasks generated by our pipeline.
-We curate a set of 100 safe website domains (different from the safety experiment, refer to Appendix E), generate task proposals using our pipeline, and measure the rate of self-reported task completion for human workers performing tasks.
-Workers start from the initial website URL in their browser and navigate pages using their mouse and keyboard while staying on the original site, reporting once the task is complete or once they believe the task is not feasible on the current site.
-We compare the feasibility rates for tasks generated by our pipeline and tasks written by human annotators given the criteria listed in Figure 2.
-Results are shown in Table 4.
-Method Feasibility Rate
+Table 1: The safety filter displays a high accuracy.
+We measure the accuracy, precision, and recall of the safety filter on a set of 100 websites, where 50 contain harmful, or mature content.
+Up to 97% of websites are correctly classified, and recall is as high as 1.0.
+Method Verifiable Rate
 Llama 3.1 70B 75%
@@
 Gemini 1.5 Pro 89%
-Human Baseline 54%
-Figure 4: Expert feasibility of proposed tasks.
-We generate tasks for 100 safe websites (listed in Appendix E), and measure the completion rates of human workers attempting to perform the generated tasks in their browser.
-Language models exceed the performance of human annotators at creating feasible web navigation tasks for agents.
-Understanding The Results.
-Language models outperform human annotators by 64.8% at creating feasible web navigation tasks.
-Larger models like Gemini 1.5 Pro display the best feasibility rates, but the smaller model Llama 3.1 70B still outperforms the human annotators by 38.9%.
-To understand the relationship between the popularity of the site being annotated and the reliability of human-written tasks, we conduct an experiment in Figure 5 comparing PageRank values (Page et al., 1999) of sites according to the official June 2024 host-level web graph from The Common Crawl Foundation (2024), versus the feasibility rates of the proposed tasks from Table 4.
-Although human annotators match the reliability of LLMs at creating feasible web navigation tasks for popular sites, LLMs outperform human annotators by 157.1% for less popular sites with low PageRank values.
-As obscurity increases, human annotators are less familiar with the sites, and the reliability of their task proposals decreases by 55.7%, whereas the reliability of tasks generated by LLMs remains relatively constant.
-This difference suggests that we should employ language models to ensure that task proposals are reliable as we begin to scale agents to vast numbers of sites on the Internet.
-However, where do we acquire this large and diverse set of websites to process for annotation?
-Figure 5: Feasibility rates vs PageRank values.
-We visualize PageRank values, a proxy for the popularity of websites, versus the expert feasibility rates of proposed web tasks.
-Human-written tasks perform on par with LLMs for popular sites, but as target sites become less popular and annotators are less familiar with them, LLMs begin to outperform human annotators at creating feasible tasks for agents.
-4.4 Scaling To 150,000 Sites
-We propose to leverage open-source crawls of the internet for large-scale task generation.
-As of June, 2024, the web graph released by The Common Crawl Foundation (2024) contains more than 300 million unique hosts, which we adapt to a data source for agents.
-In particular, we sort hosts by their PageRank values, and select the top 1M sites for task generation.
-CommonCrawl is likely to contain many sites not suitable for annotation, and the experiments in Section 4.2 illustrate that the safety filter in the task proposer can detect and remove them effectively.
-In our configuration, task generation with Llama 3.1 70B takes 14 hours for 1M sites served with vLLM (Kwon et al., 2023) on two 8-GPU nodes.
-Sections 4.2 and 4.3 show Llama 3.1 70B outperforms human annotators in safety and reliability, and we can serve it locally at significantly reduced cost compared to proprietary LLMs with a marginal loss in quality.
-The distribution of tasks generated with Llama 3.1 70B for the top 1M sites in the CommonCrawl PageRank are visualized in Figure 6.
-Understanding The Data.
-The task proposer filters out 85% of the sites in CommonCrawl, resulting in 150k sites that can be safely assigned tasks for agents.
-Visualized in Figure 6, our distribution is denser than human-written tasks, has a wide coverage of real-world sites, and diverse categories of tasks.
-We automatically label task categories (procedure in Appendix E) and find that 89% of categories have fewer than the mean of 16.9 tasks per category.
-Top categories include news search, recipe search, product lookup, tutorial search, event schedules, health information, and many more.
-Refer to Appendix E for the top categories.
-Empowered by this large and diverse collection of tasks across the internet, we can start building internet-scale agents.
-Figure 6: Distribution of 150k tasks.
-We compare the distribution of tasks generated by our pipeline (blue points) to the Mind2Web (Deng et al., 2023) dataset (orange points) via textual features extracted by a sentence embedding model, and projected in 2D with UMAP (McInnes et al., 2020).
-Our distribution is denser than human-written tasks, and has broad coverage of diverse real-world sites and tasks.
-5 Internet-Scale Agents
-In the next stage, we begin to scale LLM agents to diverse web navigation tasks across the web.
-Shown in Figure 7, we initialize a web browsing environment to the URL provided to the task proposer in Section 4, and run a language model agent to complete tasks by generating function calls in the Playwright API.
-For evaluation, current efforts typically use human-written constraints based on the final URL or page state (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; Drouin et al., 2024), but it can be difficult to scale these.
-Recall from Figure 5 that human annotators are less reliable for sites lower in the PageRank, where their familiarity is reduced.
-Results in section 4 showed that language models beat humans in safety and reliability for task generation.
-As we begin to scale agents to diverse websites from the internet, can we replace human-written criteria with language model judgments for efficient evaluation?
-Their robustness remains an important unresolved question, and current work only considers language model judges for a limited set of popular websites (He et al., 2024).
-We begin by validating the robustness of language models for evaluating diverse internet tasks.
+Table 2: Generated tasks are typically achievable.
+We measure the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer for a set of 100 websites.
+Up to 89% of tasks are achievable, and verifiable.
+Reliability is equally important for autonomous agents.
+Instructions should be followed faithfully by agents, and this requires training them with tasks that are achievable, and verifiable.
+Table 2 reports the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer in its initial phase.
+Up to 89% of tasks are achievable, and verifiable according to the study, which suggests the pipeline is producing reliable tasks.
+Together with results in Section 6, it is likely that data from the InSTA pipeline leads to agents that follow instructions faithfully.
+4.3 Scaling To 150,000 Websites
+We leverage Common Crawl for task generation.
+As of May 2025, the latest web graph released by The Common Crawl Foundation (2025) contains more than 300M unique hosts, which we adapt into a data source for agents.
+In particular, we select the top 1M sites based on their PageRank values.
+Common Crawl likely contain many unsafe websites, and these are filtered out by the task proposer.
+Each phase of task generation requires 14 hours of compute time using two 8-GPU v100 machines, and filters to 150k safe websites annotated with tasks.
+Statistics of tasks are shown in Figure 3.
+5 Internet-Scale Environment
+By this point, we have reached our first goal—to improve coverage of real-world sites by annotating 150k diverse sites with challenging agentic tasks.
+To reach our second goal, and move beyond a static dataset, towards a dynamic internet-scale environment, we require a robust evaluator for these tasks.
+Evaluation presents a subtle challenge.
+The web evolves constantly, and daily changes in website content may invalidate a fixed ground truth reference solution.
+Driven by necessity, this environment must be evaluated by a model that judges whether an agent’s solution is correct, in the latest context.
 Figure 4: Automatic evaluation for agents with language model judges.
@@
 5.1 Evaluation With Language Models
-We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved by the final timestep.
-Conditioned on a system prompt $\mathbf{y}_{\text{sys}}$, a task $\mathbf{c}$, the history of previous actions $\mathbf{a}_{1},\cdots,\mathbf{a}_{T}$ and the final observation from a web browsing environment $\mathbf{s}_{T}$ formatted in text, we generate a completion using a language model and employ a function $f^{\;\text{text}\to\text{val}}(\cdot)$ to parse the estimated probability.
-$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}\big{(}\;\text{LLM}([\;\mathbf{y}_{\text{sys}},\mathbf{c},\mathbf{a}_{1},\cdots,\mathbf{a}_{T},\text{Enc}(\mathbf{s}_{T})\;])\;\big{)}$ (4)
-Building on the sites used to measure reliability in Section 4.3, we conduct an experiment to measure the accuracy of language models to detect successful web navigation trajectories.
-We run agents on tasks generated by Llama 3.1 70B for the 100 sites in Section 4.3, and prompt language models to estimate the probability that tasks are solved $\mathbf{r}_{T}$ based on Equation 4.
-We then conduct human evaluations for the trajectories and manually assign binary success labels.
-Accuracy is calculated by applying a threshold to the predictions $\mathbf{r}_{T}>0.5$ to assign classes and tracking the rate at which the predictions agree with human labels.
-To understand robustness for sites of varying popularity, we report the accuracy of language models versus the PageRank of the corresponding sites.
-Similarly, to understand the ability of language models to judge their own uncertainty, we report their accuracy versus their prediction confidence, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$ (twice the total variation distance from the uniform distribution to the predicted distribution).
-Results are shown in Figure 8.
+We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved, and generate $\mathbf{r}_{T}$ via next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{judge}}$, a task $\mathbf{c}$, and a trajectory $\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}$.
+The LLM is instructed to produce a reasoning trace that scores the agent on a scale from 0 to 1, and embed the score as JSON in a fenced code block.
+We employ a regex template that matches to the first code block in the response, and a JSON decoder to parse $\mathbf{r}_{T}$ from the response, given by $f^{\;\text{text}\to\text{val}}(\cdot)$.
+$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}(\;\text{LLM}(\;\mathbf{x}_{\text{judge}},\mathbf{c},\text{Enc}(\mathbf{s}_{1}),\mathbf{a}_{1},\cdots,\text{Enc}(\mathbf{s}_{T}),\mathbf{a}_{T}\;)\;)$ (5)
+Verifying The Judge.
+To understand the robustness of the judge, we measure its accuracy detecting successful trajectories that were annotated by human workers.
+We annotate 100 trajectories with binary success labels, and apply a threshold $\mathbf{r}_{T}>0.5$ to obtain binary predictions from the judge.
+Figure 5 reports the accuracy of the judge as a function of the PageRank values of sites, and as a function of the confidence of the judge, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$.
+For all LLMs tested, the judge shows a high accuracy, ranging from $78.0\%$ for Gemini 1.5 Pro, to $81.7\%$ for Llama 3.1 70B, and $82.6\%$ for GPT-4o.
+Accuracy is stable as PageRank falls, suggesting the judge is accurate for less popular sites that may not be well represented in the LLM’s training data.
+Shockingly, the confidence predicted by LLMs is highly interpretable, and correlates with accuracy to the point that trajectories where conf = 1 are classified with an accuracy up to $93.1\%$.
+The emergent robustness of the judge equips us to efficiently and accurately verify agent solutions on a dynamic internet.
 Figure 5: Language models are robust evaluators.
@@
 As models become more confident, their accuracy improves (right plot), suggesting confidence is a useful proxy for the reliability of their predictions.
-Understanding The Results.
-Language models are robust evaluators for web navigation tasks.
-Accuracy remains stable relative to PageRank values, suggesting that language models are effective for sites that typical human annotators are less familiar with.
-The best results are obtained with an evaluator based on GPT-4o, which attains an accuracy of $82.6\%$, compared to $81.7\%$ for Llama 3.1 70B, and $78.0\%$ for Gemini 1.5 Pro.
-Although the accuracy is robust to PageRank, the accuracy is highly informed by confidence.
-Language models show improved accuracy as their confidence improves, suggesting that they can effectively determine when their predictions are reliable.
-When considering predictions with conf = 1, the Llama 3.1 70B evaluator displays a compelling $93.1\%$ accuracy, 0.87 precision, and 0.82 recall for detecting successful trajectories.
-Now that we can efficiently and accurately judge trajectories, we can begin to scale language model agents to diverse internet tasks and track their success.
-Harnessing this judge, we can study the current abilities and shortcomings of language model agents spanning 150k websites.
-Figure 9: Scaling LLM agents to 150k live sites.
-We run agents based on Llama 3.1 70B to complete tasks generated by our pipeline.
-We estimate success probabilities using a language model evaluator (left plot), and estimate probabilities agents are on the right track (right plot).
-16.7% of tasks are estimated to be successful with conf = 1, and the spread of probabilities suggests data spans many difficulties.
+Figure 6: Statistics for our agent reasoning dataset.
+We conduct a large data collection experiment using our top checkpoint for Qwen3 1.7B.
+Our dataset has 2.2M screenshots, 2.2M reasoning traces for actions, and 150k traces for the judge.
+50.0% of the trajectories are successful according to the judge, and have diverse ratings for efficiency and self-correction.
 5.2 Scaling To 150,000 Agents
-We scale language model agents to 150k live sites in diverse domains across the Internet and attempt to complete 150k web navigation tasks generated by our pipeline.
-Shown in Figure 9, we evaluate the trajectories using a Llama 3.1 70B judge and run agents based on Llama 3.1 70B, selected because this model demonstrates high accuracy in Figure 8, and running currently available propriety models would be prohibitively expensive at this scale (see Appendix K for a cost analysis with different LLMs).
-We find that agents solve 16.7% of tasks with a model confidence of conf = 1.
-Furthermore, we observe that 35k tasks are judged to be on the right track with a confidence of conf = 1, suggesting that these could be solved if a larger computing budget was allocated.
-The spread along the x-axis in both plots in Figure 9 suggests that our tasks cover a broad range of difficulties, and working to solve them presents an opportunity to improve the capabilities of LLM agents.
-We observe that, when judging success, our evaluator tends to prefer binary predictions with high confidence values, suggesting that this subset of predictions is accurate based on Figure 8 results.
-Analyses for the agents that produced Figure 9 are presented in Appendix F.
-Assembling the stages of the InSTA pipeline, we first generated tasks for 1M diverse sites across the Internet, we then dispatched language model agents to complete tasks for 150k sites marked as safe, and finally we evaluated the trajectories using a language model judge.
-Equipped with this large and diverse source of data, we seek to understand the utility of the data for training agents.
+With our task proposer, agent, and judge driven by pretrained language models, we have all components needed to harness internet-scale data.
+We conduct a large data collection experiment, running language model agents to complete tasks on 150k websites from our official dataset, producing 2.2M screenshots and 2.2M reasoning traces for actions within 150k trajectories.
+The judge annotates these trajectories, producing 150k reasoning traces for evaluations, and leading to the statistics in Figure 6.
+For this experiment, we employ a fine-tuned Qwen3 1.7B as the agent (refer to the next section), and Qwen3 235B zero-shot as the judge.
+Data collection requires 1,200 v100 GPU hours, and costs \$521.55 based on current AWS spot instance pricing, a fraction of the budget that industry labs are spending towards agents.
+If you have the right data, this small budget is sufficient and no human annotations are required to build models that compete with frontier LLMs as agents.
 6 Training Agents
-We compare the performance of agents trained on data from the InSTA pipeline to agents trained on human demonstrations from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), two recent and popular benchmarks for web navigation.
-Recent works that mix synthetic data with real data control the real data sampling probability in the batch $p_{\text{real}}$ independently from data size (Trabucco et al., 2024).
-We employ $p_{\text{real}}=0.5$ in few-shot experiments and $p_{\text{real}}=0.8$ otherwise.
-Shown in Figure 9, our data have a wide spread in performance, so we apply several filtering rules to select high-quality training data.
-First, we require the evaluator to return conf = 1 that the task was successfully completed, and that the agent was on the right track (this selects data where the actions are reliable, and directly caused the task to be solved).
-Second, we filter data where the trajectory contains at least three actions.
-Third, we remove data where the agent encountered any type of server error, was presented with a captcha, or was blocked at any point.
-These steps produce $7,463$ high-quality demonstrations in which agents successfully completed tasks on diverse websites.
-We sample 500 demonstrations uniformly at random from this pool to create a diverse test set, and employ the remaining $6,963$ demonstrations to train agents on a mix of real and synthetic data.
-6.1 Improving Data-Efficiency
-Figure 10: Data from InSTA improves efficiency.
-Language model agents trained on mixtures of our data and human demonstrations scale faster than agents trained on human data.
-In a setting with 32 human actions, adding our data improves Step Accuracy by +89.5% relative to human data for Mind2Web, and +122.1% relative to human data for WebLINX.
-In a data-limited setting derived from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), agents trained on our data scale faster with increasing data size than human data alone.
-Without requiring laborious human annotations, the data produced by our pipeline leads to improvements on Mind2Web that range from +89.5% in Step Accuracy (the rate at which the correct element is selected and the correct action is performed on that element) with 32 human actions, to +77.5% with 64 human actions, +13.8% with 128 human actions, and +12.1% with 256 human actions.
-For WebLINX, our data improves by +122.1% with 32 human actions, and +24.6% with 64 human actions, and +6.2% for 128 human actions.
-Adding our data is comparable in performance gained to doubling the amount of human data from 32 to 64 actions.
-Performance on the original test sets for Mind2Web and WebLINX appears to saturate as the amount of human data increases, but these benchmark only test agent capabilities for a limited set of 150 popular sites.
-6.2 Improving Generalization
-Figure 11: Our data improves generalization.
+Figure 7: InSTA unlocks great potential in small models.
+We train agents based on Qwen 3 1.7B using trajectories produced by a Qwen 3 235B data collection policy, and optionally filtered by a Qwen 3 235B judge (see Judge Filtered vs.
+Uniformly Sampled).
+We report success rates on a test set of 3,000 held-out websites and tasks.
+Before training, Qwen 3 1.7B has a zero-shot success rate of 11.5% according to a Qwen 3 235B judge, and we improve this by +45.3% absolute percentage points.
+Our top checkpoint outperforms the Qwen 3 235B data collection policy, and Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller.
+Notably, filtering with a Qwen 3 235B judge leads to agents that improve according to independent secondary judges, including Gemini 2.5 Flash, Llama 4 Maverick, and Gpt 4.1 Nano, suggesting it generalizes well.
+We’ve built an internet-scale data flywheel that produces trajectories annotated with scores from a judge that can help us train LLM agents.
+To understand the quality of data produced by the flywheel, we conduct a series of experiments training models on the data, and testing on popular benchmarks.
+These experiments focus on three main questions: (1) what is the impact of increasing data scale?
+(2) do agents transfer to new domains?
+(3) do agents scale with test-time compute?
+6.1 Performance Improves With Data Scale
+The recurring lesson in deep learning is that large-scale high-quality data wins, but researchers are struggling to materialize this promise for agents (Xue et al., 2025).
+Our paper aims to solve the data problem blocking researchers from materializing this promise, and this experiment provides a valuable signal that scaling high-quality data allows small models to compete with strong LLMs from top industry labs.
+To proceed, we collect 20k trajectories using a Qwen 3 235B data collection policy, annotated with scores from a Qwen 3 235B judge.
+We then train agents based on Qwen 3 1.7B with SFT on varying scales of the data.
+Results in Figure 7 show that performance improves with increasing data scale, and gains scale faster on data filtered by the judge.
+To filter the data, we select trajectories where Judge(Success) = 1.
+Our top checkpoint outperforms the Qwen 3 235B data collection policy, and beats Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller.
+The trend in the figure suggests there is room to scale further.
+Figure 8: Our agents zero-shot transfer to WebVoyager.
+With no additional training or specialized data, our checkpoints for Qwen 3 1.7B in Section 6.1 zero-shot transfer to the WebVoyager benchmark.
+Trends found on our test set appear to hold for WebVoyager as well, and our top checkpoints for Qwen 3 1.7B continue to match frontier LLMs in performance for three of four judges.
+6.2 Agents Transfer To New Domains
+Statistically correct evaluation for deep learning models requires a test dataset that does not overlap with the training dataset, but researchers do not agree on how to implement this guideline for agents.
+Recent works train agents on the same websites they test on (Murty et al., 2025; Zhou et al., 2024c; Putta et al., 2024), which may obfuscate progress (Xue et al., 2025).
+Our next experiment shows that we can implement a stronger train-test split.
+Agents trained with our data can zero-shot transfer to WebVoyager (He et al., 2024) without any data from the benchmark.
+Results in Figure 8 show our Qwen 3 1.7B checkpoint matching strong LLMs on WebVoyager (He et al., 2024) for three of four judges, confirming trends in Section 6.1.
+Our ability to zero-shot transfer relatively small agents to WebVoyager (He et al., 2024) makes it likely that our pipeline leads to capable agents.
+Figure 9: Our data transfers to static benchmarks.
 We train agents with all human data from the WebLINX and Mind2Web training sets, and resulting agents struggle to generalize to more diverse test data.
 Adding our data improves generalization by +149.0% for WebLINX, and +156.3% for Mind2Web.
-To understand how agents trained on data from our pipeline generalize to diverse real-world sites, we construct a more diverse test set than Mind2Web and WebLINX using 500 held-out demonstrations produced by our pipeline.
-Shown in Figure 11, we train agents using all human data in the training sets for WebLINX and Mind2Web, and compare the performance with agents trained on 80% human data, and 20% data from our pipeline.
-Agents trained with our data achieve comparable performance to agents trained purely on human data on the official test sets for the WebLINX and Mind2Web benchmarks, suggesting that when enough human data are available, synthetic data may not be necessary.
-However, when evaluated in a more diverse test set that includes 500 sites not considered by existing benchmarks, agents trained purely on existing human data struggle to generalize.
-Training with our data improves generalization to these sites by +149.0% for WebLINX agents, and +156.3% for Mind2Web agents, with the largest gains in generalization Step Accuracy appearing for harder tasks.
+Static Benchmarks.
+To complement the online evaluation in Figure 8, we also explore how our data impacts agents trained on static benchmarks.
+We first train baseline agents on human demonstrations from the Mind2Web (Deng et al., 2023) and WebLINX (Lù et al., 2024) datasets.
+We write a preprocessor to convert our data into the expected action-observation formats these benchmarks use (which involves discarding our reasoning trace).
+With data converted, we compare the baseline to agents trained on a mix of 80% human data, and 20% our data, and test on (1) their official test set, and (2) 500 diverse sites from our official test set.
+Results in Figure 9 show that agents trained with our data perform equally well on the original test sets for these benchmark, but generalize better to our harder test set.
+Overall, we see +149.0% for WebLINX agents, +156.3% for Mind2Web agents, and gains in Step Accuracy on our test set are larger for the harder tasks.
+Additional experimental details are listed in Appendix H.
+On three popular benchmarks (WebVoyager, WebLINX, Mind2Web), our pipeline leads to capable agents and does not depend on human annotations.
+6.3 Performance Scales With Test-time Compute
+To understand the ability for web agents to scale with additional test-time compute, we ablate the number of tokens in the reasoning budget.
+Figure 10 shows the success rate as a function of the reasoning budget for Gemini 2.5 Flash, the top-performing agent we tested.
+There is a monotonic improvement in the success rate as the reasoning budget increases, and the trend suggests that performance may not be saturated with 500 reasoning tokens.
+Training language model agents to reason before taking actions is a promising path to better agents, and we are releasing a large reasoning dataset for multimodal agents to study this.
+Our dataset contains 2.2M screenshots, 2.2M reasoning traces for actions, 150k traces for judge evaluations, and led to the results in Section 6.1.
+The data will be linked on our website, alongside an official huggingface dataset for tasks.
+Figure 10: Agents improve with a larger reasoning budget.
+We ablate the number of tokens in the reasoning budget for the top-performing agent, and see a monotonic improvement in the success rate as the reasoning budget increases.
+Gemini 2.5 Flash has a 70% success rate with a budget of 500 reasoning tokens, up from 60% for a budget of 50 tokens.
+The scaling of performance with the reasoning budget highlights a promising behavior in successful web agents that can be studied.
 7 Conclusion
-We present a pipeline that unlocks new scaling potential for language model agents by generating internet-scale web navigation data for diverse sites.
-We employ pretrained language models to generate, attempt, and evaluate web navigation tasks at scales beyond what humans can manually annotate.
-Our results show that language models outperform human annotators at detecting safe websites, and proposing feasible web navigation tasks.
-In addition, we find that language models can judge successful web navigation trajectories with an accuracy up to 93.1% for their most confident predictions.
-We scale task generation to the top 1M sites in the CommonCrawl PageRank, and test agents on 150k live web navigation tasks.
-Agents based on Llama 3.1 70B solve 16.7% of these tasks zero-shot.
-In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data.
-When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
-Our work reveals several exciting directions for future work.
-First, our work can be scaled further.
-The latest CommonCrawl release contains data for more than 300 million sites, suggesting another 1,000 times more data than we explored could be available for training agents by scaling the pipeline.
-In addition, our language model judge was employed offline, and its high accuracy suggests that it could be used to guide an online algorithm.
-Finally, we considered primarily text-based agents in this work, and our pipeline could be extended to generate data for multimodal tasks.
+In the spirit of deep learning, we have developed an approach to efficiently harness internet data for LLM agents, and have unlocked a dynamic internet-scale environment.
+In building this environment, we presented a method to annotate 150k diverse sites with challenging agentic tasks, and showed how training on data from our pipeline allows small models to compete with frontier LLMs as agents, on a fraction of the budget.
+Our pipeline consists of a task proposer, agent, and judge driven by pretrained language models that together curate high-quality data for agents, without human intervention.
+Our top checkpoint for Qwen 3 1.7B has a success rate of 56.9% on our test environment, outperforming the data collection policy Qwen 3 235B, beating the 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash, while being smaller and faster than these.
+Our models zero-shot transfer to WebVoyager, and scale with test-time compute.
+We are releasing the entire pipeline, including code, models and data, so that it may serve as a foundation for researchers to build the next generation of language model agents with internet data.
+7.1 Future Work
+Our work reveals several exciting directions in future work.
+First, our work can be scaled further: the latest Common Crawl release contains data for more than 300M sites, suggesting another 1,000 times more data could be available for agents by scaling the pipeline.
+In addition, we trained agents to optimize the judge scores indirectly via filtered SFT, and the judge’s high accuracy suggests that it could be optimized via reinforcement learning instead.
+RL is especially promising for how it can improve reasoning capabilities in agents.
+Finally, while the data we collect is multimodal, we focus on textual tasks in this paper, and our pipeline could be extended to produce multimodal tasks.
 8 Acknowledgments
@@
 We also thank faculty who provided feedback and direction for parts of this work, including Aviral Kumar, Graham Neubig, Shuyan Zhou, and others.
+Finally, we thank the open source community for developing systems and methods for serving LLMs efficiently, without which this paper would not be possible.
 Appendix A Limitations & Safeguards
@@
 For instance, agents visiting shopping sites can influence the statistics produced by analytics tools, which can impact prices on products, and product decisions from companies.
-Furthermore, agents visiting harmful content can add such harmful content to datasets, and perpetuate harmful behaviors into the training data for future agents.
-We mitigate these risks by carefully designing the task proposal stage of the InSTA pipeline.
+Furthermore, agents seeing harmful content on the web can add that content to datasets inadvertently, and propagate harmful behaviors to future agents.
+We mitigate these risks in the design of the task proposal stage.
 We consider the risks posed to analytics tools by limiting the engagement between agents and sites.
-We generate only one task per website, and we limit agents to just 10 actions per site, which includes clicks, typing, dropdown selection actions, and more.
-By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $30$ seconds of interaction per site on average).
+We generate only one task per website, and we limit agents to just 30 actions per site, which includes clicks, typing, dropdown selection, and more.
+By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $90$ seconds of interaction per site on average).
 By utilizing data from the InSTA pipeline in an offline fashion, as in Section 6 of the main paper, no additional web traffic is generated when training agents.
 To ensure that agents do not modify the state of the web (i.e.
-avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide an instruction in the system prompt of the task proposer (see Figure 2) to avoid writing tasks that require the agent to modify the state of the web.
-The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate (such as social media, and forums).
+avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide instruct the task proposer (see Figure 2) to avoid writing tasks that require the agent to interact with personal data, or user accounts.
+The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate, including social media, and forum sites.
+There is likely a manner to safely train agents to operate user accounts, but we leave this task to future researchers.
 We explore the performance of the task proposer at filtering out unsuitable sites in Section 4.2, and find that all models detect unsuitable sites with a recall from $0.98$ to $1.0$, and accuracy up to 97%, suggesting our filter is reliable.
-Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and include harmful, and mature content.
+Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and thoroughly test the safety filter.
+To remove Personally Identifiable Information (PII) from the data used for training agents, we include scrubadub, an industry standard PII removal tool for python developed by Leap Beyond, a data consultancy based in the European Union.
+Our pipeline has an argument that toggles the usage of scrubadub to remove PII from all website data, and we recommend this option be set.
 Appendix B Ethical Considerations
-One important ethical consideration when gathering data from the internet is to handle copyrighted, private, and sensitive materials carefully.
-The internet contains vast amounts of personal data created by users that includes personally-identifying-information that should not be included in public datasets.
-We address this ethical consideration in two ways.
-First, the task proposer is instructed to filter out social media sites and forums that are likely to contains personally-identifying-information.
-Second, we store and release only the prompts we used, and traces for agents’ actions—importantly, we do not release any web source code that could be used to recover sensitive data.
-These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are included in our data, and methods for detecting, replacing, and removing such materials from datasets remains an important task for researchers working on safety.
+One important ethical consideration when harnessing internet data is to carefully handle copyrighted, private, and sensitive materials.
+The internet contains vast amounts of PII, which should be avoided when training models.
+We address this in two ways.
+First, we instruct the task proposer to filter out sites that may contain PII, including social media websites, and forums.
+Second, our pipeline has an argument that toggles the usage of scrubadub to remove PII, and we recommend this option be set.
+To prevent copyrighted materials from being used for training, we do not allow the agent to copy any books, documents, videos, audio, or files.
+The information present in the HTML DOM is publically available, and is converted into a sparse Markdown summary that focuses on interactive elements, and significantly transforms the underlying webpage in a manner that is not reversible.
+These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are shown to the agent, and methods for detecting, replacing, and removing such materials from text and images remains an important task for researchers working on safety.
 Appendix C Broader Impacts
-As their capabilities broaden, language model agents are being increasingly used to operate real-world systems and APIs.
+As their capabilities broaden, language models are being used to interface with real-world systems.
 This shift comes with several benefits and risks.
 Agents that operate your computer to aid in work tasks can significantly boost productivity for certain workers, but can displace others whose jobs have been fully automated.
-Agents that operate web browsers to complete personal tasks for users can provide convenience, but expose a new attack vector where compromised agents perform unintended actions.
-Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential in nature, and harder to address purely from a research perspective.
-Our data pipeline aims to facilitate internet-scale training for agents using offline data.
-Training from offline data is important because current agent capabilities are low, and training them on live websites risks taking bandwidth away from real users, while polluting the internet with noisy LLM outputs.
-Due to these concerns, we advocate for researchers building internet-scale agents to follow the guidelines set forth in this work: (1) agents should not modify the state of the internet by making accounts, creating posts, or submitting content; (2) agents should be throttled to a maximum rate of interaction, and a maximum number of interactions per website; (3) copyrighted, private, and sensitive materials should be removed from training data.
+Agents that operate browsers to complete personal tasks provide convenience, but expose a new attack vector where rogue agents perform unintended actions.
+Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential, harder to address purely with new agent research, and may require policy changes.
 Appendix D Agents.txt & Standards For Internet Agents
-Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites—if at all.
+Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites, and what information on webpages agents are allowed to see.
 Desireable controls include rate limits for interactions, limits for maximum numbers of interactions, restrictions to allow agents to interact with certain pages and not others, and restrictions on the kind of data on webpages that agents are allowed to observe (achieved via tagging elements to hide their content from agents).
-In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from their primary user-facing platform.
+In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from user-facing pages.
 Appendix E More Details On Task Generation
@@
 For each of the following domains, provide a realistic, and specific task that a user could reasonably accomplish in a single session on the website, and limit your response to 20 words.
-Figure 12: System prompt for task generation.
-We carefully design the system prompt for task generation to ensure that sites not suitable for inclusion in the training data for agents are detected and removed.
-This prompt ensures that proposed tasks are passive in nature, and only involve retrieving information—active tasks like making posts and creating accounts are explicitly not allowed.
-We provide the system prompt used for task generation in Figure 12.
+Figure 11: System prompt for the exploration phase of task generation.
+We design the system prompt for task generation to detect and remove unsafe websites.
+This prompt ensures that tasks are passive, and do not modify content on a website.
+Refer to the next figures for the in-context examples used for the task proposer, and the system prompt used in the feedback step.
+We provide the system prompt used in the first phase of the task generation loop in Figure 11.
 This prompt was provided to Llama 3.1 70B, GPT-4o, and Gemini 1.5 Pro to generate tasks and filter sites unsuitable for annotation in Section 4.
-We carefully designed this system prompt to enforce that generated tasks are passive in nature, and do not modify the state of content on the internet.
+We carefully designed this system prompt to enforce that generated tasks are passive, and do not modify content on a website.
 In addition to this system prompt, we employed a list of 100 hand-picked in-context examples of website URLs and appropriate tasks, which are provided in the following JSON list.
-When querying an LLM to generate tasks, we randomly sample 16 in-context examples from the following list, and provide only these examples to the LLM.
-This helps promote diverse tasks.
+When querying an LLM, we randomly sample 16 in-context examples from the list, and provide only these examples to the LLM to generate a task to guide exploration of the site.
+This improves diversity in the exploration phase.
 "domain": "archive.org",
@@
 "task": "Locate the most recent update on the global biodiversity status."
+We also provide the system prompt used in the second phase of the task generation loop, where trajectories from agents are fed back to the task proposer, which generates a harder, grounded task.
+This prompt instructs the task proposer to create a challenging task based on how an expert user could be expected to use the shown website.
+The task proposer also predicts a list on intermediate steps that can be used as a hint for agents, and a success criteria that can be used to improve the verifier.
+You are a helpful assistant designing tasks for a web automation script.
+I will show you previous runs of the script, including previous tasks, webpages, actions, and performance reviews, formatted in markdown.
+Help me design *challenging* new tasks.
+## Formatting The Proposed Task
+Format your task in the following JSON schema:
+‘‘‘json
+"proposed_task": str,
+"steps": List[str],
+"criteria": str
+Here is what each key means:
+- proposed_task‘: A specific, challenging task that an expert user might leverage this website to complete.
+- Must not require making an account, logging in, submitting personal information, making a purchase, or placing an order.
+- steps‘: Steps an expert user would follow to complete the proposed task.
+- criteria‘: The required answer, and criteria to determine if the task was completed.
+## Example Tasks For Inspiration
+Suppose you want to design a task around the C-to-C Hose-Shut-Off Valve on awg-fittings.com’:
+‘‘‘json
+"proposed_task": "What is the C-to-C Hose-Shut-Off Valve length in mm?",
+"steps": [
+"Navigate to awg-fittings.com’",
+"Open the product catelog for fittings",
+"Locate the product listing for the C-to-C Hose-Shut-Off Valve",
+"Find the product length in mm, and respond with that length in the answer"
+"criteria": "The answer should include the specific length of ’237 mm for this product"
+Suppose you want to design a task around the document The Angora cat; how to breed train and keep it on biodiversitylibrary.org’:
+‘‘‘json
+"proposed_task": "Open a scanned copy of The Angora cat; how to breed train and keep it’.",
+"steps": [
+"Navigate to biodiversitylibrary.org’",
+"Search for The Angora cat; how to breed train and keep it in the search bar",
+"Click on the title of the document in the search results",
+"Confirm the correct document is displayed in an embedded PDF reader"
+"criteria": "The final webpage should display the correct document in an embedded PDF reader"
+Suppose you want to design a task around the Generative Adversarial Networks paper on scholar.google.com’:
+‘‘‘json
+"proposed_task": "How many citations does the paper Generative Adversarial Networks have?",
+"steps": [
+"Navigate to scholar.google.com’",
+"Search for Generative Adversarial Networks in the search bar",
+"Locate the correct paper in the search results",
+"Find an up-to-date citation count, and respond with that count in the answer"
+"criteria": "The answer should include an up-to-date citation count, which is ’80613’ as of April 2025"
+Suppose you want to design a task around the word serendipity on wiktionary.org’:
+‘‘‘json
+"proposed_task": "What is the definition and etymology of the word serendipity’?",
+"steps": [
+"Navigate to wiktionary.org’",
+"Search for serendipity in the search bar",
+"Find the definition and etymology sections of the serendipity page",
+"Summarize the contents of these sections in the answer"
+"criteria": "The answer should mention Serendip (or Serendib), coined by English writer and politician Horace Walpole in 1754"
+Thanks for helping me design challenging new tasks, please follow the instructions carefully.
+Start your response with an analysis for how an expert user would leverage this website, followed by a step-by-step breakdown of your proposed task, and finally, enter your task in the JSON format.
+Respond in 500 words.
 E.1 Details For Safety Experiment
-This list of examples is also provided in our code release, alongside the script that we used to generate task proposals for the top 1M sites in the CommonCrawl PageRank (The Common Crawl Foundation, 2024).
-Using these prompts for task generation, we can filter our sites that are unsuitable for annotation, due to containing harmful content, or sensitive user data.
+Using these prompts for task generation, we remove unsafe websites.
 To evaluate the performance of our filter, we employed a set of 100 curated websites, where 50 are manually verified as safe, and 50 are manually verified as unsafe based on the filtering conditions.
-These sites were chosen to span popular sites that typical annotators are likely familiar with, and less popular sites that annotators may not already be familiar with.
+These sites were chosen to span popular sites that typical users are likely familiar with, and less popular websites that may be underrepresented in LLM training data.
 safe_sites_list = [’dhss.mo.gov’, dizionari.corriere.it’, southgippsland.vic.gov.au’, ds.iris.edu’, lobbycontrol.de’, ’4rsmokehouse.com’, barnsleyfc.co.uk’, wiwi.uni-wuerzburg.de’, uplandca.gov’, lsus.edu’, wpcode.com’, webopedia.internet.com’, tamko.com’, premierchristian.news’, genome.jgi.doe.gov’, burgerking.ca’, thehugoawards.org’, radio.fm’, thevinyldistrict.com’, unilang.org’, raywhitegroup.com’, grapevinetexas.gov’, sanfrancisco.cbslocal.com’, hyde-design.co.uk’, breastcancerfoundation.org.nz’, ludwigsburg.de’, ignitionrobotics.org’, deliverit.com.au’, kodokan.org’, clickstay.com’, searchdatamanagement.techtarget.com’, oceanario.pt’, wentworthpuzzles.com’, catholicworldreport.com’, quizlet.com’, innovation.nhs.uk’, synonyms.reverso.net’, news.siemens.co.uk’, readability-score.com’, co.modoc.ca.us’, cityofmyrtlebeach.com’, loire.gouv.fr’, lawphil.net’, saem.org’, parmigianoreggiano.it’, engaging-data.com’, itf-tkd.org’, aka.education.gov.uk’, ub.uni-kl.de’, mottchildren.org’]
 unsafe_sites_list = [’refer to the code release’]
-We provide the list of safe sites in the preceding code block, and in the code release, but we provide the list of unsafe sites in just the code release to keep this manuscript clean of harmful content.
+We provide the list of unsafe sites in the code to keep this manuscript clean.
 Using these lists, we conduct an experiment measuring the accuracy, precision, and recall of our safety filter for detecting sites that are not suitable for training agents.
-In this experiment, we run the task proposer with the same prompts used to scale annotation the top 1M sites in the CommonCrawl PageRank, and we consider a site to be marked positive for unsafe content if the task proposer generates “N/A” rather than predicting a task.
-To calculate the human baseline for this experiment, we showed participants the system prompt observed by the task proposer, and asked participants to label sites as safe or unsafe using this per-example prompt:
-[…]
-Human participants were not allowed to visit the URL shown, and had to determine whether the site is safe for annotation purely from their prior knowledge (the same conditions faced by the task proposer).
-The 100 sites for the safety experiment were shuffled into a uniformly random order to ensure the order of annotation did not bias the human annotators predictions.
-One human participant was used to obtain the human baseline result in Table 3.
+In this experiment, we run the initial exploration phase of the task proposer with the same prompts used in the main experiments (shown in the last section), and we consider a website to be marked positive for unsafe content if the task proposer generates “N/A” rather than a task.
+We then calculate accuracy, precision, and recall for various LLMs.
 E.2 Details For Reliability Experiments
-Similar to the previous safety experiment, we employed human participants to obtain a human baseline for task feasibility.
-In particular, we showed human participants the system prompt in Figure 12 for the task proposer, and had them write a task for each of the following websites without visiting the URL (the same conditions faced by the task proposer).
-The following 100 sites were shuffled into a uniformly random order to ensure the participants were not influenced by the order in which sites were shown.
-After tasks were proposed by participants, and by LLMs, we evaluated the expert feasibility of tasks by manually attempting to complete the tasks proposed by each set of participants, and marking tasks as feasible, or not feasible based on our own ability to complete them.
-In total, we annotated 400 tasks, which required 8 hours of annotation.
-One human participant was used to obtain the human baseline result in Table 4.
+We evaluated the verifiable rate of tasks generated by the initial phase of task generation manually.
+For a set of 100 randomly sampled websites marked as safe by the task proposer in the initial phase, we attempted to complete the generated task, and checked that enough information is present on the website that a solution can be verified.
+In total, we annotated 300 tasks for Table 2 in 6 hours, and provide the 100 website URLs used in this experiment in the following code block.
 reliability_sites_list = [’godaddy.com’, chrome.google.com’, apple.com’, support.cloudflare.com’, support.apple.com’, edition.cnn.com’, go.microsoft.com’, google.de’, w3.org’, yandex.ru’, bfdi.bund.de’, microsoft.com’, apps.apple.com’, networksolutions.com’, support.mozilla.org’, yelp.com’, cnn.com’, ec.europa.eu’, developer.mozilla.org’, icann.org’, books.google.com’, globenewswire.com’, onlinelibrary.wiley.com’, gnu.org’, slideshare.net’, metacpan.org’, porkbun.com’, oag.ca.gov’, spiegel.de’, linuxfoundation.org’, help.opera.com’, mayoclinic.org’, podcasts.apple.com’, nhs.uk’, addons.mozilla.org’, google.fr’, pewresearch.org’, finance.yahoo.com’, weforum.org’, g2.com’, savethechildren.org’, news.com.au’, biblia.com’, yr.no’, engadget.com’, microsoftstore.com’, ema.europa.eu’, theintercept.com’, princeton.edu’, foodandwine.com’, sfgate.com’, voguebusiness.com’, ourworldindata.org’, livingwage.org.uk’, cms.law’, msdmanuals.com’, websitesetup.org’, support.xbox.com’, treehugger.com’, tripadvisor.com.pe’, mondragon.edu’, greenparty.ca’, aaojournal.org’, restaurantpassion.com’, iwillteachyoutoberich.com’, moneyconvert.net’, gesundheitsinformation.de’, ovc.uoguelph.ca’, zdnet.be’, oxfordamerican.org’, snackandbakery.com’, journals.uic.edu’, confused.com’, standards.globalspec.com’, onlyinyourstate.com’, ahsgardening.org’, wyze.com’, nornickel.ru’, viessmann.fr’, benetton.com’, firecomm.gov.mb.ca’, executedtoday.com’, eukn.eu’, fraeylemaborg.nl’, verizon.com/about/news-center’, orthodoxalbania.org’, cheapjoes.com’, bake-eat-repeat.com’, plattformpatientensicherheit.at’, hifinews.com’, cellsignal.com’, thenotariessociety.org.uk’, chosenfoods.com’, westerndressageassociation.org’, pridesource.com’, northtacomapediatricdental.com’, strade-bianche.it’, pvdairport.com’, institute.sandiegozoo.org’, raintaxi.com’]
-[…]
 E.3 Automatic Task Categorization
+We employ Llama 3.1 70B to categorize tasks.
+We prompt Llama 3.1 70B with the system prompt in Figure 12 to assign a category in 3 words or less to encourage simple categories.
+Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean.
+There is occasional overlap between categories, which can be observed in Figure 13, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated.
+We provide our task categorization script in the official code release.
 You are a helpful scientific assistant categorizing tasks on the web.
@@
 Using these categories, we seek to understand agent performance by category.
-To better understand the statistics of generated tasks, we employ Llama 3.1 70B to assign task categories.
-We prompt Llama 3.1 70B with the system prompt in Figure 13 to assign a category in 3 words or less to encourage simple categories.
-Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean.
-There is occasional overlap between categories, which can be observed in Figure 14, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated.
-We provide our task categorization script in the official code release.
 Appendix F Understanding Agent Capabilities & Limitations
-Figure 14: Largest categories for internet-scale task generation.
-We assign categories to 150k web navigation tasks generated by our pipeline in Section 4, and visualize the number of tasks for each of the largest 70 categories.
-Top categories include article search, news search, recipe search, product lookup, and more.
+Figure 13: Largest categories for task generation.
+We categorize 150k tasks generated by our pipeline in Section 4, and visualize the number of tasks in the largest 70 categories.
+Top categories include article search, news search, recipe search, and product lookup.
 The top 12 task categories have more than 1600 tasks assigned to each of them, the mean number of tasks per category is 16.9, and 89% of categories (7741 in total) have fewer than the mean number of tasks.
-Figure 15: Most successful categories for internet-scale task generation.
-We explore the rates of task completion for the top categories of tasks generated by our pipeline.
-We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 of such categories.
-Results show that 22 categories are solved with more than a 50% rate with agents based on Llama 3.1 70B.
-To complete the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently.
+Figure 14: Most solved categories for task generation.
+We explore the completion rates for the top categories of tasks generated by our pipeline.
+We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 categories.
+Results show that 22 of these categories are solved with more than a 50% rate with zero-shot agents based on Llama 3.1 70B.
+To complement the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently.
 Shown in Figure 14, we plot the average judge success probability prediction $\mathbf{r}_{T}$ versus task category for the top 70 most successful categories that have at least 100 tasks assigned to them.
-Based on the figure, top categories include search for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research.
-Based on these results, the top 22 categories are solved with more than a 50% rate using agents based on Llama 3.1 70B running zero-shot.
+Based on the figure, top categories include searching for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research.
+Based on these results, the top 22 categories are solved with an average success probability $>0.5$ using zero-shot agents based on Llama 3.1 70B.
 As stronger models are developed, the success rates for agents running in our pipeline are likely to improve, and the quality of the data we generate will jointly improve.
 In addition to studying the best-performing categories, we also explore the limitations of current agents via their least successful categories.
-Shown in Figure 16, we select the bottom 70 categories in terms of their average judged success probability for categories with at least 100 tasks assigned to them.
+Shown in Figure 15, we select the bottom 70 categories via their average judge success probability for categories with at least 100 tasks assigned.
 Many of these categories require agents to remember and reason about previous interactions, such as the product comparison category.
-For this category, an agent must review several products, and compare their details from memory.
-In these cases, access to a note-taking app may improve performance.
-Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening and playing files.
-While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment could unlock these abilities in future work.
+For this category, an agent must review several products, and compare details from memory.
+In these cases, access to a note-taking tool may improve performance.
+Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening files.
+While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment with applications pre-installed could unlock these abilities in future work.
 Figure 15: Least successful categories for internet-scale task generation.
@@
 Appendix G Agent & Judge System Prompts
-We provide the system prompt used with our agent below.
-This prompt is released in our official code, alongside the observation processor that maps webpage DOM to a compact markdown format, referenced in the system prompt.
-You are a helpful assistant operating my web browser.
-I will show you webpages formatted in markdown, and I want your help to complete a web navigation task.
-Read the webpage, and respond with an action in JSON to interact with the page, and help me complete the task.
-## Formatting The Response
-Respond with actions in the following JSON schema:
+We provide the system prompt that powers the agent in this paper.
+This prompt is released in the code, alongside a fast HTML to Markdown processor we built.
+The agent prompt is carefully designed to elicit reasoning capabilities, and the experiment in Figure 10 shows the prompt is successful.
+You are helping me complete tasks by operating a web browser. I will share the current task, and a sequence of webpages and actions from previous steps.
+## Action Instructions
+Based on the information we discovered so far, and the progress we made in previous steps, you are helping me determine the next action.
+You will provide an action as JSON in a fenced code block:
 ‘‘‘json
@@
 "action_kwargs": dict,
-"target_element_id": int
-Here is what each key means:
-- action_key‘: The action to perform.
-- action_kwargs‘: Named arguments for the action.
-- target_element_id‘: The id of the element to perform the action on.
-## Available Actions
-Im using playwright, a browser automation library, to interact with the page.
-Im parsing the value assigned to action_key into a method call on the page object, or an element object specified by the value assigned to target_element_id‘.
-Here are the available actions:
+"target_element_id": int | null
+Actions have the following components:
+- action_key‘: The name of the selected action.
+- action_kwargs‘: A dictionary of arguments for the action.
+- target_element_id‘: An optional id for the element to call the action on.
+## Action Definitions
+Ive prepared an API documentation below that defines the actions we can use to complete the task.
 ### Click Action Definition
@@
 ### Example Click Action
-Suppose you want to click the link ‘[id: 5] Sales link‘:
+Suppose you want to click ‘[id: 5] Sales link‘:
 ‘‘‘json
@@
 ### Example Hover Action
-Suppose you want to hover over the image ‘[id: 2] Company Logo image‘:
+Suppose you want to hover over ‘[id: 2] Company Logo image‘:
 ‘‘‘json
@@
 "target_element_id": 2
+### Scroll Action Definition
+- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
+- delta_x‘: The number of pixels to scroll to the right.
+- delta_y‘: The number of pixels to scroll down.
+### Example Scroll Action
+Suppose you want to scroll down the page by 300 pixels:
+‘‘‘json
+"action_key": "scroll",
+"action_kwargs": {
+"delta_x": 0,
+"delta_y": 300
+"target_element_id": null
 ### Fill Action Definition
@@
 - value‘: The text value to fill into the element.
-### Example Fill Action
-Suppose you want to fill the input ‘[id: 13] "Name..." (Enter your name text field)‘ with the text John Doe‘:
+### Example Fill Action (Text Input)
+Suppose you want to fill ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
 ‘‘‘json
@@
 "target_element_id": 13
+### Example Fill Action (Range Slider)
+Suppose you want to set ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘.
+The slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘.
+You must translate the desired \$1000 to the correct underlying value of ‘20‘:
+‘‘‘json
+"action_key": "fill",
+"action_kwargs": {
+"value": "20"
+"target_element_id": 71
 ### Select Action Definition
@@
 ### Example Select Action
-Suppose you want to select the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
+Suppose you want to select the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
 ‘‘‘json
@@
 "target_element_id": 67
+### Set Checked Action Definition
+- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
+- checked‘: Boolean value to check or uncheck the checkbox.
+### Example Set Checked Action
+Suppose you want to check ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
+‘‘‘json
+"action_key": "set_checked",
+"action_kwargs": {
+"checked": true
+"target_element_id": 21
 ### Go Back Action Definition
@@
 ### Example Go Back Action
+Suppose you want to go back to the previous page:
 ‘‘‘json
@@
 ### Example Goto Action
-Suppose you want to open google search:
+Suppose you want to open the DuckDuckGo search engine:
 ‘‘‘json
@@
 "action_kwargs": {
-"url": "https://www.google.com"
+"url": "https://www.duckduckgo.com"
 "target_element_id": null
 ### Stop Action Definition
-- stop‘: Stop the browser when the task is complete, or the answer is known.
-- answer‘: Optional answer if I requested one.
+- stop‘: Stop when the task is complete, and report your progress.
+- answer‘: Optional answer sent back to me.
 ### Example Stop Action
+Suppose the task is complete, and you want to stop and report your progress:
 ‘‘‘json
@@
 "action_kwargs": {
-"answer": "Im done!"
+"answer": "The desired task is now complete."
 "target_element_id": null
-Thanks for helping me perform tasks on the web, please follow the instructions carefully. Start your response with an explanation in 50 words, and choose exactly one action you would like to perform.
-We also provide the system prompt used with out LLM judge.
-The system prompt instructs the judge to predict a json-formatted dictionary that contains a “success” key, and an “on_right_track” that represent the estimated probability that the task is successful, and that the agent is on the right track towards solving the task, respectively.
-These distinctions are adapted from Koh et al.
-(2024b), and help us filter for high-quality training data by distinguishing trajectories that were solved by the agent’s own actions from trajectories that were solved by chance.
-You are a helpful assistant providing feedback on a web automation script.
-I will show you a list of previous actions, the current webpage formatted in markdown, and the proposed next action.
-I want your help evaluating the proposed action, to determine if the desired task is complete, or if we are on the right track towards future completion.
-## Reading The Action Schema
-You will see actions in the following JSON schema:
+## Formatting Your Response
+Write a 200 word revised plan based on new information we discovered, and progress we made in previous steps. After your response, provide the next action as JSON in a fenced code block.
+We also provide the system prompt used by the judge.
+The system prompt instructs the judge to predict JSON within a fenced code block that contains a “success” key, an “efficiency” key, and a “self_correction” key.
+The success key represents a score from 0 to 1 that estimates the probability the task is successfully completed.
+The efficiency key represents a score from 0 to 1 that estimates the probability the agent has taken the most efficient path to solve the task.
+The self correction key represents a score from 0 to 1 that estimates the probability that the agent has demonstrated self corrective behaviors during its completion of the task.
+These behaviors include when the agent backtracks to a more promising state, re-plans when new information is discovered relevant to the task, and recognizes its own mistakes.
+These are generally behaviors we expect from successful agents, but for this paper we only filter by the success key to select training data for agents.
+You are helping me evaluate a browser automation script. I will share a task provided to the script, and a sequence of webpages and actions produced by the script.
+## The Action Format
+The script produces actions as JSON in a fenced code block:
 ‘‘‘json
@@
 "target_element_id": int
-Here is what each key means:
-- action_key‘: The action to perform.
-- action_kwargs‘: Dictionary of arguments for action.
-- target_element_id‘: The id of the element to perform the action on.
-## Available Actions
-Im using playwright, a browser automation library, to interact with the page.
-Im parsing the value assigned to action_key into a method call on the page object, or an element specified by the value assigned to target_element_id‘.
-Here is an example action:
+Actions have the following components:
+- action_key‘: The name of the selected action.
+- action_kwargs‘: A dictionary of arguments for the action.
+- target_element_id‘: An optional id for the element to call the action on.
+## Action Definitions
+Ive prepared an API documentation below that defines the actions the script can use to complete the task.
+### Click Action Definition
+- click‘: Click on an element specified by target_element_id‘.
 ### Example Click Action
-Here is an example where the script clicked the link ‘[id: 5] Sales link‘:
+Here is an example where the script clicks ‘[id: 5] Sales link‘:
 ‘‘‘json
@@
 "target_element_id": 5
+### Hover Action Definition
+- hover‘: Hover over an element specified by target_element_id
+### Example Hover Action
+Here is an example where the script hovers over ‘[id: 2] Company Logo image‘:
+‘‘‘json
+"action_key": "hover",
+"action_kwargs": {},
+"target_element_id": 2
+### Scroll Action Definition
+- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
+- delta_x‘: The number of pixels to scroll to the right.
+- delta_y‘: The number of pixels to scroll down.
+### Example Scroll Action
+Here is an example where the script scrolls down the page by 300 pixels:
+‘‘‘json
+"action_key": "scroll",
+"action_kwargs": {
+"delta_x": 0,
+"delta_y": 300
+"target_element_id": null
+### Fill Action Definition
+- fill‘: Fill an input element specified by target_element_id with text.
+- value‘: The text value to fill into the element.
+### Example Fill Action (Text Input)
+Here is an example where the script fills ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
+‘‘‘json
+"action_key": "fill",
+"action_kwargs": {
+"value": "John Doe"
+"target_element_id": 13
+### Example Fill Action (Range Slider)
+Here is an example where the script sets ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘.
+This slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘.
+The script translates the desired \$1000 to the correct underlying value of ‘20‘:
+‘‘‘json
+"action_key": "fill",
+"action_kwargs": {
+"value": "20"
+"target_element_id": 71
+### Select Action Definition
+- select‘: Select from a dropdown element specified by target_element_id‘.
+- label‘: The option name to select in the element.
 ### Example Select Action
-Here is an example where the script selected the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
+Here is an example where the script selects the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
 ‘‘‘json
@@
 "target_element_id": 67
+### Set Checked Action Definition
+- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
+- checked‘: Boolean value to check or uncheck the checkbox.
+### Example Set Checked Action
+Here is an example where the script checks ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
+‘‘‘json
+"action_key": "set_checked",
+"action_kwargs": {
+"checked": true
+"target_element_id": 21
+### Go Back Action Definition
+- go_back‘: Go back to the previous page (‘target_element_id must be null).
+### Example Go Back Action
+Here is an example where the script goes back to the previous page:
+‘‘‘json
+"action_key": "go_back",
+"action_kwargs": {},
+"target_element_id": null
+### Goto Action Definition
+- goto‘: Navigate to a new page (‘target_element_id must be null).
+- url‘: The URL of the page to navigate to.
 ### Example Goto Action
-Here is an example where the script opened google search:
+Here is an example where the script opens DuckDuckGo search:
 ‘‘‘json
@@
 "action_kwargs": {
-"url": "https://www.google.com"
+"url": "https://www.duckduckgo.com"
 "target_element_id": null
+### Stop Action Definition
+- stop‘: Stop when the task is complete, and report the progress.
+- answer‘: Optional answer from the script.
 ### Example Stop Action
-Here is an example where the script stopped with the message "Im done!":
+Here is an example where the script stops and reports its progress:
 ‘‘‘json
@@
 "action_kwargs": {
-"answer": "Im done!"
+"answer": "The desired task is now complete."
 "target_element_id": null
-## Formatting The Response
-Think step by step, and start your response with an explanation of your reasoning in 50 words. Then, provide an evaluation in the following JSON schema:
+## Evaluation Instructions
+Based on the progress of the script, you are helping me determine if the desired task has been completed successfully.
+You will provide scores as JSON in a fenced code block:
 ‘‘‘json
 "success": float,
-"on_right_track": float,
-Here is what each key means:
-- success‘: What is the probability the desired task has been completed successfully, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
-- on_right_track‘: What is the probability the script is on the right track towards a future success, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
-Thanks for helping me evaluate the script, please follow the instructions carefully. Start your response with a step by step explanation. Then, provide an evaluation in the JSON schema above.
+"efficiency": float,
+"self_correction": float
+### Score Definitions
+- success‘: Your confidence the desired task has been completed successfully.
+- range: 0.0 (not possible) to 1.0 (absolutely certain).
+- efficiency‘: Your confidence the script has taken the most efficient path to complete the task.
+- range: 0.0 (not possible) to 1.0 (absolutely certain).
+- self_correction‘: Your confidence the script has demonstrated self-corrective behaviors during its completion of the task.
+These behaviors include backtracking to a more promising state, replanning when new information is discovered, and recognizing its own mistakes.
+- range: 0.0 (not possible) to 1.0 (absolutely certain).
+Write a 300 word analysis that establishes specific criteria to rigorously evaluate whether the task was completed, followed by which criteria the script has satisfied.
+After your response, provide your scores as JSON in a fenced code block.
 Appendix H Details For Training Agents
-To understanding the utility of the generated data for training agents, we filter the data, and compare our filtered data to human demonstrations on the WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023) benchmarks.
-In particular, we sweep over different sizes of random subsets of human actions, from 32 to 256, which helps us understand the value of synthetic data generated from the InSTA pipeline versus different scales of human data.
-We then fine-tune models based on google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks.
-We employ identical training hyperparameters to those used in Lù et al.
-(2024) and Deng et al.
-(2023) to ensure that our results are comparable to previous work.
-Section H reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites.
-In order to prepare our data, we employ three filtering rules.
-In the first rule, we filter for data where the agent was predicted to have succeeded at the task with conf = 1, and was predicted to be on the right track with conf = 1.
-This filtering rule is motivated by our findings in Section 5, where we found that our LLM judge based on Llama 3.1 70B has an accuracy up to 93.1% at detecting successful trajectories for its predictions with conf = 1.
-Filtering based on both “success” and “on_right_track” conditions is essential to obtain data where the agent directly caused the task to be solved, rather than the task being solved by external conditions.
-The next filtering rule we use is to select trajectories with at least three actions, which helps create training data that is not too easy (i.e.
-not solved after just one or two actions).
-Finally, we select tasks where the agent did not encounter any errors during execution.
-These include being presented with server errors such as 404 Not Found, and 403 Forbidden, encountering a captcha, and being blocked, even if just temporary, from the target website.
-These filtering steps produce an automatically curated set of $7,463$ demonstrations from our pipeline where agents successfully completed tasks generated by the InSTA pipeline.
-We reserve 500 demonstrations from this pool for our test set, and the rest for training agents in Figure 10.
-The original Mind2Web dataset contains $2,350$ tasks.
-Appendix I Additional Related Works
-While writing this paper, concurrent work was released that introduces a Proposer-Agent-Evaluator framework for web navigation agents (Zhou et al., 2024c).
-There are several key differences between our work and theirs, and the most important difference is scale.
-We generate tasks for 1M sites on the internet, whereas their work considers just 85 real-world sites, 5 sites from WebArena (Zhou et al., 2024b), and 13 sites from WebVoyager (He et al., 2024).
-The second difference is evaluation.
-Safety and reliability play crucial roles when gathering data, and we conduct an analysis on the safety and reliability of data generated by our method on 100 real-world sites.
-Another major difference pertains to offline learning.
-Offline learning should be used when scaling agents because current agent capabilities are low, and training them online risks polluting the internet with noisy LLM outputs, while taking bandwidth away from real users.
-The final difference pertains to the train-test split.
-We train agents on diverse internet data, and transfer to target benchmarks, while the agents presented in Zhou et al.
-(2024c) train on sites from target benchmarks using synthetic tasks.
-Our train-test split is stronger, and evaluates the ability for agents trained on our synthetic data to generalize to novel websites, domains, and tasks.
-Appendix J Hyperparameters
+To understand the utility of data we obtained, we train agents and test on four relevant benchmarks: InSTA, WebVoyager (He et al., 2024), Mind2Web (Deng et al., 2023), WebLINX (Lù et al., 2024).
+In particular, our test set consists of a held-out set of 3,000 websites and tasks produced by the task generation feedback loop.
+Note these websites are not present in the training set.
+For WebVoyager (He et al., 2024), we transfer agents trained on our data zero-shot to 643 tasks on 15 websites WebVoyager (He et al., 2024).
+The websites in the WebVoyager benchmark are not present in the 20k trajectories we collected in Section 6.1 we used for training agents.
+For this experiment, we fine-tuned models based on Qwen 3 1.7B with a maximum sequence length of 16,384 tokens, and the most recent 5 observations, and actions in the context.
+We employed full fine-tuning on this model, with Adam, learning rate of 5e-5, batch size of 32, bfloat16, and other parameters kept as the PyTorch defaults for Adam.
+Each model was trained using one epoch, a linear warm-up corresponding to the first 0.01 steps of training, and a linear decay to 6e-5 afterwards.
+To filter data, we select trajectories that were scores as Judge(Success) = 1, which corresponded to 10.5k of 20k trajectories produced by the Qwen 3 235B data collection policy.
+Scores for filtering were produced by a Qwen 3 235B judge.
+We employed a simple filtering strategy that only considers the success score from the judge, and no other filtering conditions were used.
+Note the judge also predicts efficiency, and self correction scores, which could likely also help select the best data for training, but we did not explore filtering by these scores in this work.
+For experiments on static benchmarks, we fine-tune google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks.
+We employ identical training hyperparameters to those used by Lù et al.
+(2024) for Llama in their official training code and Deng et al.
+(2023) for Flan to ensure that our results are comparable to previous work.
+Section 6.2 reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites.
+The websites in these static benchmarks were not present in the dataset produced by InSTA for this experiment, to ensure fairness.
+Appendix I Hyperparameters
 We provide a list of the hyperparameters used in this work in Table 3.
-Values are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to employ standard hyperparameters for training agents on WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023).
-On the Mind2Web benchmark, we did not observe a loss in performance for training longer on few-shot experiments, and kept training iterations fixed between few-shot and all-available-data experiments.
-This is likely due to the FLAN-T5 model used in Mind2Web being too small to overfit to the data.
-For experiments on WebLINX, we found that performance for models trained on just human data degrades if models are trained for the full $10,000$ steps, so a smaller $2,000$ steps was chosen.
+Hyperparameters for mixing our data with human data on static benchmarks are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to adhere to standard hyperparameters for WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023).
+We train using all available human data on these benchmarks, and add trajectories filtered using the previously discussed methodology, sampled at a 20% rate in the data-loader compared to an 80% rate for human data.
 Hyperparameter Name Value
-OpenAI Model Name gpt-4o
-Google Model Name gemini-1.5-pro
-Llama Model Name meta-llama/Llama-3.1-70B-Instruct
-CommonCrawl PageRank Revision cc-main-2024-apr-may-jun-host-ranks.txt.gz
+Models Used For Agents Qwen/Qwen3-1.7B
+Qwen/Qwen3-235B-A22B
+meta-llama/Llama-4-Maverick-17B-128E-Instruct
+meta-llama/Llama-3.1-70B-Instruct
+meta-llama/Llama-3.3-70B-Instruct
+google/gemini-2.5-flash
+Models Used For Judges Qwen/Qwen3-235B-A22B
+meta-llama/Llama-4-Maverick-17B-128E-Instruct
+meta-llama/Llama-3.1-70B-Instruct
+meta-llama/Llama-3.3-70B-Instruct
+google/gemini-2.5-flash
+google/gemini-1.5-pro
+openai/gpt-4.1-mini
+openai/gpt-4o
+Common Crawl PageRank cc-main-2024-apr-may-jun-host-ranks.txt.gz
 Number of sites before filtering $1,000,000$
 Number of tasks after filtering $146,746$
-Max Tokens Per Observation $4,096$
-Max Tokens Per Action $2,048$
-Max Tokens Per Judgement $2,048$
-Max Tokens Per Task $64$
-Max Observations Per Agent Context 5
-Max Actions Per Agent Context 5
-Max Observations Per Judge Context 1
-Max Actions Per Judge Context 5
-OpenAI Inference API Sampling Temperature 0.5
-OpenAI Inference API Sampling Top P 1.0
-Mind2Web HuggingFace Model Name google/flan-t5-large
-Mind2Web Few-Shot Training Iterations $11,505$
+Max Tokens Per Observation $2,048$
+Max Tokens Per Agent Trace $1,024$
+Max Tokens Per Judge Trace $1,024$
+Max Tokens Per Task Proposer Trace $1,024$
+Last Steps Per Agent Context 5
+Last Steps Per Judge Context 5
+Last Steps Per Task Proposer Context 5
+Task Proposer Feedback Loops 1
+LLM Sampling Temperature 0.5
+LLM Sampling Top P 1.0
+LLM Sampling Top K default
+Fine-tuned LLM in Section 6.1 Qwen/Qwen3-1.7B
+InSTA Training Epochs 1
+InSTA Batch Size $32$
+InSTA Learning Rate 5e-5
+InSTA Optimizer Adam
+Mind2Web LLM google/flan-t5-large
 Mind2Web Training Iterations $11,505$
@@
 Mind2Web Learning Rate 5e-5
-WebLINX HuggingFace Model Name meta-llama/Llama-3.1-8B-Instruct
-WebLINX Few-Shot Training Iterations $2,000$
-WebLINX All Training Iterations $10,000$
+Mind2Web Optimizer Adam
+WebLINX LLM meta-llama/Llama-3.1-8B-Instruct
+WebLINX Training Iterations $10,000$
 WebLINX Batch Size $16$
 WebLINX Learning Rate 5e-5
-Filtering Success Threshold 1.0
-Filtering On Right Track Threshold 1.0
-Few-Shot $p_{\text{real}}$ 50%
-All Human Data $p_{\text{real}}$ 80%
+WebLINX Optimizer Adam
+Data Filtering Condition Judge(Success) = 1
+Human Data Sampling Probability $p_{\text{real}}$ 80%
 Table 3: Hyperparameters used in our paper.
-We organize hyperparameters into five sections, including names of language model backbones, parameters of the data generation pipeline, sampling parameters for the OpenAI inference API, training parameters used by corresponding benchmarks, and filtering parameters used to prepare our data for training agents in Section 6.
-Appendix K Cost Analysis For Llama 3.1 70B
-To contextualize how Llama 3.1 70B is important for a project at this scale, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models.
-As the analysis shows, using Llama 3.1 70B is most feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise in accuracy and performance.
-We have deep gratitude for the Llama team at Meta working to make developments in language modeling available to the research community at no cost.
+We organize hyperparameters into seven sections, for the names of LLMs used as agents in the paper, the names of LLMs used as judges in the paper, the hyperparameters used for data collection, the sampling parameters for LLMs, the training parameters for static benchmarks, and the filtering and data mixing hyperparameters.
+Appendix J Cost Analysis For Llama 3.1 70B
+To understand the significant reduction in cost that we obtain by running LLMs locally to generate data, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models.
+As the analysis shows, using Llama 3.1 70B is a feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise performance.
+We have deep gratitude to the Llama team at Meta, and the Qwen team at Alibaba for working to make developments in language modeling available to the research community at no cost.
+We see up to a 95% reduction in cost with these models.
 Variable Name Value
 Number of tasks $146,746$
-Typical tokens per observation $1,024$
+Average tokens per observation $1,024$
 Max observations per agent context window $5$
-Typical agent / judge response size $128$
+Average agent / judge response size $512$
 Max tokens per system prompt $1,024$
-Max steps per task $10$
-Estimated tokens processed by the agent
-$146,746*(((1,024+128)*5+1,024)*10-(1,024+128)*15)=$ $7,419,477,760$
-Tokens processed by the judge
-$146,746*(1,024+1,024+128*10)=$ $488,370,688$
-Total tokens processed $7,907,848,448$
-Expected API cost for GPT-4o \$ $39,539.24$
-Expected API cost for Gemini 1.5 Pro \$ $9,489.42$
+Average steps per task $15$
+Estimated tokens processed by the agent 14.65B tokens
+Tokens processed by the judge 1.35B tokens
+Total tokens processed 16.00B tokens
+Expected API cost for GPT-4.1 \$ 32,000.00
+Expected API cost for Gemini 2.5 Pro \$ 20,000.00
 Expected AWS compute cost for serving Llama 3.1 70B
-       (14 days for two 8-gpu v100 spot instances) \$ $6,622.56$
-Percent saved using Llama 3.1 70B $[30.2,83.3]$
+       (3,840 v100 GPU hours using spot instances) \$ $1,575.70$
+Percent saved using Llama 3.1 70B $[95.08,92.12]$
 Table 4: Cost analysis for different LLM models in the fully-scaled pipeline.